### v1

# NP-Hardness and ETH-Based Inapproximability of Communication Complexity via Relaxed Interlacing

Abstract
We prove that computing the deterministic communication complexity $D\left(f\right)$ of a Boolean function is NP-hard for any number of alternation rounds, a problem posed by Yao [5] when introducing communication complexity. In order to build our reduction, we expand on a suite of structural lemmas about interlaced constructions that were used by Mackenzie and Saffidine [4] to refute the strong form of the direct sum conjecture. These can be re‑used as black boxes in future lower‑bound constructions.
Only constant alternations required. The optimal protocol on the instances our reduction maps onto require only constant round alternation. This proves NP-hardness for unrestricted rounds of alternation on a more restrictive class of functions than those with optimal protocols requiring a growing number of rounds of communication. This also separates our result from recent concurrent and independent work by Hirahara et al. [3], whose hardness proof relies on unbounded alternations.
Recursive embedding and gap amplification. The self-similar nature of our gadgets allows for recursive embedding. We leverage this to provide a proof sketch showing that, under the ETH, our construction yields an additive inapproximability gap of $\omega(1)$. We further outline how this technique could be extended to prove NP-hardness of approximating $D\left(f\right)$ within a constant additive error. A full, rigorous treatment of these inapproximability results is deferred to a future version of this work.
Flexibility of the framework. Our NP-hardness result showcases the power of the modular framework for amplifying communication lower bounds introduced in [4] and further developed in this paper. The generic structural theorems we present give access to whole class of other constructions/reductions which warrant further investigation. Beyond settling the complexity of communication complexity, we expect it to be of use in the broader understanding of communication complexity.
footnotetext: Concurrent and independent work by Hirahara, Ilango, and Loff [3] also establishes NP-hardness of deciding $D(f)\geq k$, but their reduction requires an unbounded number of alternations and does not extend to ETH-based additive inapproximability.
1 Preliminaries
1.1 Notation & Parameters
Table 1 gives a list of important notation used throughout the paper. Table LABEL:tab:master-params gives a list of parameters and their dependencies throughout the paper.
Symbol Meaning / Comment
$\mathbb{N}$ $\{0,1,2,\dots\}$, the set of non‑negative integers
$[k]$ Shorthand for $\{1,\dots,k\}$
$\left\lceil x\right\rceil_{\!2}$ the least power of two that is at least $x$
R,CR,\;C Generic finite row / column index sets
$\mathsf{Rows}(M)$ Row index set of a Boolean matrix $M$
$\mathsf{Cols}(M)$ Column index set of $M$
$M\!\upharpoonright_{R^{\prime}\times C^{\prime}}\;\;$ (or $\varepsilon\left(M;R^{\prime},C^{\prime}\right)$) Restriction (“extraction”) of $M$ to rows $R^{\prime}$ and columns $C^{\prime}$
$M^{\top}$ Transpose of $M$
$f\circ g$ Binary interlacing of $f$ and $g$ (Def. 1)
$f\left\langle k\right\rangle$ kk‑fold interlace of a function $f$ (Def. 2)
$\widehat{R}=[k]\times R$ Interlaced row index set produced by $\cdot\left\langle k\right\rangle$
$S_{q,t}(C)$ (q,t)(q,t)‑balanced column multiset (Def. 9)
$f\left\langle q,S_{q,t}(C)\right\rangle$ Relaxed interlace, restricted to $S_{q,t}(C)$ domain
$M\left[p,x,y\right]$ Family of pp‑fold interlaced sub‑matrices (Def. 8)
(Q,p)(Q,p)‑equipartition Row subset with $\lvert U_{q}\rvert\!\geq p$ for each $q\!\in\!Q$ (Def. 5)
[…]
$\varepsilon_{q,t}$ Generic small accuracy / statistical‑distance parameter
$D\left(M\right)$ Deterministic communication complexity of $M$
$\Delta,\;\delta,\;\;b,$ Robustness parameters (Def. 11)
$t$ Power‑of‑two parameter controlling #of interlaced games
B,rB,\;r Protocol budget and extension factor (Extension Thm.)
$\phi,\mathcal{M}_{1},\mathcal{M}_{2},\dots$ Canonical matrices in the reduction chain (Sec. 4)
Table 1: Frequently used notation. Unless stated otherwise, all logarithms are base 22.
Table 2: Consolidated list of parameters, their ranges, and where they first arise. All inequalities are understood to hold simultaneously whenever the corresponding lemma/theorem is invoked.
Symbol Meaning Constraints Usage Notes
$b$ robustness depth; sets row fraction $2^{-b}$ $b\geq 1$ Def. 11 Upper‑bounds $t$ via $t\leq 2^{b}$
$\delta$ baseline column margin $0<\delta\leq\frac{1}{\sqrt{2}}-\tfrac{1}{2}$ Def. 11 Fixed for whole paper
$\varepsilon_{q,t}$ AGHP error of (q,t)(q,t)‑balanced set $\varepsilon_{q,t}:=(qt|C|)^{-c}$ §2 Drives $y^{\prime}$ in Lem. 10
$q$ total outer blocks $q\geq t$; $q=rt$ in Lem. 14 Def. 10
$t$ # of fully independent blocks $1\leq t\leq 2^{b}$; power of two; $\Delta\leq\log t$ Lem. 10 Lem. 13 Inner interlace factor
$r$ replication with $q=rt$ $r\geq 1$ (int.) Lem. 14 Adds $\lceil\log r\rceil$ to budget $B$
$x$ row density for each $p$ interlaced functions in $\cdot\left[p,x,y\right]$ $0<x\leq 1$; often $x=2^{-b}$ or $2^{-B}$ Def. 8 Must satisfy $x\leq 2^{-k}$ in Lem. 4
$y$ column density $0<y\leq 1$; e.g. $\frac{1}{2}+\delta$ Def. 8 Raised to powers in projection lemmas
$\Delta$ parameter controlling number of interlace operations required to get $\delta$ $1\leq\Delta\leq\log t$ Lem. 13 Defines $p=2^{\Delta-1}+1$
$B$ protocol bit budget $B=D\left(M\right)+\log t+\lceil\log r\rceil$ (Extension Theorem). Forces protocol to spend all block bits
$h,h^{\prime}$ column‑density maps $(1+\varepsilon)h^{\prime}(t)2^{-B}r+\varepsilon\geq h(t)$ (Extension Theorem). Otherwise monotone
$B^{\prime}$ One round budget in Lem. 14 $B=\log t+\lceil\log r\rceil$ Lem. 14 Any optimal protocol must use all $B$ bits
1.2 Definitions
Definition 1 (Binary Interlacing Operation).
Let $f\colon X_{1}\times Y_{1}\to\{0,1\}$ and $g\colon X_{2}\times Y_{2}\to\{0,1\}$ be functions. Define
$X^{\prime}=X_{1}\sqcup X_{2}\quad\text{and}\quad Y^{\prime}=Y_{1}\times Y_{2}.$
Then the function $f\circ g\colon X^{\prime}\times Y^{\prime}\to\{0,1\}$ is defined by
$(f\circ g)(x,y)=\begin{cases}f(j,y_{1})&\text{if }x=(1,j)\text{ and }y=(y_{1},y_{2}),\\[2.84526pt] g(j,y_{2})&\text{if }x=(2,j)\text{ and }y=(y_{1},y_{2}).\end{cases}$
Definition 2 (Interlacing Operation).
Let $f:X\times Y\to\{0,1\}$ be a function, $k$ an positive integer. Define the sets
$X^{\prime}\coloneqq\bigsqcup_{i\in[k]}X\quad\text{and}\quad Y^{\prime}\coloneqq Y^{k}.$
We then define the function $f\left\langle k\right\rangle:X^{\prime}\times Y^{\prime}\to\{0,1\}$ by setting
$f\left\langle k\right\rangle(x,y)$ $=f(j,y_{i})$ $\text{with }x\!=\!(i,j)\in X^{\prime}\text{ and }y\!=\!(y_{1},\ldots,y_{k})\!\in\!Y^{\prime}\text{.}$
Matrices as set-indexed functions.
For finite sets $R$ (rows) and $C$ (columns), a matrix is a function $M:R\times C\to\{0,1\}$. We write $\mathsf{Rows}(M)=R$ and $\mathsf{Cols}(M)=C$.
Definition 3 (Extraction Operator).
Let $M:R\times C\to\{0,1\}$ and let $R^{\prime}\subseteq R$, $C^{\prime}\subseteq C$. The extraction (restriction) of $M$ to $R^{\prime}$ and $C^{\prime}$ is
$\varepsilon\left(M;R^{\prime},C^{\prime}\right):=M\!\upharpoonright_{R^{\prime}\times C^{\prime}}.$
Definition 4 (pp‑Equipartitioned Row Set).
Fix a finite set $R$ and an integer $k\geq 1$. Let the interlaced row index set be $\widehat{R}:=[k]\times R$.
For an integer $p\geq 1$ and a subset $U\subseteq\widehat{R}$ write
$U_{i}\;:=\;\{\,j\in R:(i,j)\in U\,\}\quad\text{for each }i\in[k].$
We say $U$ is a $p$-equipartition of $\widehat{R}$ (with respect to the blocks $[k]$) if
$\lvert U_{i}\rvert\;\geq\;p\quad\text{for every }i\in[k].$
If equality $\lvert U_{i}\rvert=p$ holds for every $i$, we call $U$ an exact $p$-equipartition (of $\widehat{R}$).
Definition 5 ((Q,p)(Q,p)‑Equipartitioned Row Set).
Fix a finite set $R$, an integer $k\geq 1$ and subset $Q\subseteq[k]$. Let the interlaced row index set be $\widehat{R}:=[k]\times R$.
For an integer $p\geq 1$ and a subset $U\subseteq\widehat{R}$ write
$U_{q}\;:=\;\{\,j\in R:(q,j)\in U\,\}\quad\text{for each }q\in Q.$
We say $U$ is a $(Q,p)$-equipartition of $\widehat{R}$ if
$\lvert U_{q}\rvert\;\geq\;p\quad\text{for every }q\in Q.$
If equality $\lvert U_{q}\rvert=p$ holds for every $q$, we call $U$ an exact $(Q,p)$-equipartition (of $\widehat{R}$).
Definition 6.
For functions $\phi$ and $\phi^{\prime}$, we say $\phi^{\prime}$ is a subgame of $\phi$, denoted $\phi^{\prime}\sqsubseteq\phi$, if we can select a number of rows and columns from the matrix representing $\phi$ and rearrange their intersection into $\phi^{\prime}$. A supergame is the dual notion of a subgame.
We can extend the subgame relation to sets of functions. Conceptually a set of functions $A$ is a subgame of a set of functions $B$ if all games in $B$ have a subgame in $A$.
Definition 7.
For sets of functions $\Phi$ and $\Phi^{\prime}$, we say that a set $\Phi^{\prime}$ is a subgame of set $\Phi$ if all matrices in $\Phi$ have a subgame in $\Phi^{\prime}$. That is $\Phi^{\prime}\sqsubseteq\Phi$ iff $\forall\phi\in\Phi,\exists\phi^{\prime}\in\Phi^{\prime},\phi^{\prime}\sqsubseteq\phi$.
Definition 8 (Bracket Notation).
Given an $m\times n$ matrix $M$, a positive integer $p$, and real numbers $0<x,y\leq 1$, define
$T=\lceil mx\rceil.$
Then we define the set
$M\left[p,x,y\right]\;=\;\left\{\varepsilon\left(M\left\langle p\right\rangle;R,C\right)\;\middle|\;\begin{array}[]{l}R\subseteq[p]\times\mathsf{Rows}(M)\text{ is an exact }T\text{-equipartition},\\[2.58334pt] C\subseteq\mathsf{Cols}(M),\;\;|C|=\lceil n^{p}y\rceil\end{array}\right\}.$
1.3 Lemmas on the Interlace Operation
In this section we introduce lemmas proved in Mackenzie and Saffidine [4] regarding the interlace operation.
Lemma 1 (Monotonicity Lemma).
Given an $m$ by $n$ matrix $M$, for any, $p^{\prime}\leq p$, $0<x^{\prime}\leq x\leq 1$, and $0<y^{\prime}\leq y\leq 1$, we have
$D\left(M\left[p^{\prime},x^{\prime},y^{\prime}\right]\right)\leq D\left(M\left[p,x,y\right]\right).$
Lemma 2 (Extended Maximum Projection Lemma).
$\forall 1\leq l\leq p,D\left(M\left[p,x,y\right]\right)\geq D\left(M\left[l,x,y^{\frac{l}{p}}\right]\right)$.
Lemma 3 (Partition Lemma).
For $M$, $p\in\mathbb{N}$, $p\geq 1$, $0\leq\tau\leq 1$, $\delta\in\{0,1\}$, $0<x\leq\frac{1}{2}$, $0<y\leq 1$, if $D\left(M\left[2p,2x,y/4\right]\right)\geq 1$, then
$D\left(M\left[2p+\delta,2x,y\right]\right)$ $\geq 1+D\left(M\left[p+\delta,x,y\right]\right)$ (1)
$D\left(M\left[2p,2x,y\right]\right)$ $\geq 1$
$\hskip 5.69046pt+\!\min\left(\begin{aligned} &D\left(M\left[p,x,y^{\frac{1}{1+\tau}}\right]\right),\\ &D\left(M\left[p(1-\tau)+1,x,y^{\frac{\tau}{1+\tau}}\right]\right)\end{aligned}\right).$ (2)
Lemma 4.
For any matrix $M$, integers $s\leq k$, $p\in\mathbb{N}$, reals $0<x\leq\frac{1}{2^{k}}$, $0<y\leq 1$, $2<\rho$, if $D\left(M\left[p,x,y/4\right]\right)\geq 1$ and $(\rho-1)^{k}\leq\rho^{k-s}$, then
$D\left(M\left[2^{k}\left(\frac{\rho-1}{\rho-2}\right)^{s}p,2^{k}x,y^{\rho^{s}}\right]\right)\geq k+D\left(M\left[p,x,y\right]\right).$
Lemma 5 (Specific Hard Seed Lemma).
For $\phi=\{0,1\}$, $p\in\mathbb{N}$ and $0<x,y\leq 1$, we have
$D\left(\phi\left[p,x,y\right]\right)\geq\left\lceil\log(p+\log y)\right\rceil+1.$
Lemma 6.
For any $0<x,y\leq 1$, and integers $k,m\geq 0$, we have
$k+m+D\left(\phi_{i}\left[p,x,y\right]\right)\geq D\left(\phi_{i}\left[p,\min\{1,2^{k}x\},\min\{1,2^{m}y\}\right]\right)$.
Lemma 7.
$D\left(M\left[1,x,y\right]\right)=D\left(M^{\top}\left[1,y,x\right]\right)$.
2 Relaxing the Interlace Operation
To keep the domain of (Interlacing Operation). polynomial while retaining its key projection properties, we introduce a relaxed interlacing operation that is $k$-wise independent only up to a carefully chosen accuracy. Rather than leaving this accuracy as a free parameter, we hard-wire a value that is more than sufficient for all later proofs.
2.1 Almost k-Wise Independent Sets
Fixed parameters.
Throughout this subsection we fix
  1. i)
    an integer $q\geq 1$ (number of interlacing layers),
  2. ii)
    a balance threshold $t\geq 1$ (number of columns we control),
  3. iii)
    an alphabet size $p\geq 2$ for the column set $C=\{0,\dots,p-1\}$.
We then define a deterministic accuracy level
$\varepsilon_{q,t}:=(qt|C|)^{-c}$
Bit-level representation.
Let $\ell=\lceil\log p\rceil$ and fix an injective encoding $\mathsf{enc}:C\to\{0,1\}^{\ell}$ (the final codeword may be unused). A tuple $\mathbf{c}=(c_{1},\dots,c_{q})\in C^{q}$ is mapped to the bit string $\mathsf{enc}(\mathbf{c})\in\{0,1\}^{q\ell}$ obtained by concatenating the block encodings.
Almost $k$-wise independent support in $\{0,1\}^{q\ell}$.
With $k:=t\ell$, the construction of Alon et al. [1] yields a multiset [^1] A set would suffice, but the original analysis is stated for multisets. $\mathcal{B}\subseteq\{0,1\}^{q\ell}$ of size
||= $\mathcal{B}|\;=\;2^{\,O\!\left(\log\log(q\ell)+k+\log(1/\varepsilon_{q,t})\right)}\;=\;(qp)^{\,O(t)}\cdot\operatorname{polylog}(q),$ (3)
such that the uniform distribution on $\mathcal{B}$ is $\varepsilon_{q,t}$-close (in statistical distance) to $k$-wise independence. Importantly, this set can be constructed in $(qp)^{\,O(t)}\cdot\operatorname{polylog}(q)$ time.
Definition 9 (Balanced column domain).
Define the multiset
$\mathcal{S}\;:=\;\left\{\,\mathsf{enc}^{-1}(\mathbf{b})\;\bigm{|}\;\mathbf{b}\in\mathcal{B}\right\}\;\subseteq\;C^{q},$
where $\mathsf{enc}^{-1}$ is applied block-wise, dropping any tuple that decodes to the unused codeword $p$. We call $\mathcal{S}$ the $(q,t)$-balanced column set.
Balanced-projection guarantee.
Projecting onto any subset of at most $t$ column indices examines at most $k=t\ell$ of the underlying bits, so the AGHP guarantee transfers directly:
Lemma 8 (Balance under $t$ projections).
For every index set $J\subseteq[q]$ with $|J|\leq t$ the marginal of $\mathcal{S}$ on $J$ is $\varepsilon_{q,t}$-close to the uniform distribution on $C^{J}$; equivalently, every pattern in $C^{J}$ appears with frequency
$\frac{1}{p^{|J|}}\;\pm\;\varepsilon_{q,t}.$ (4)
Proof.
The $\ell|J|\leq k$ bits encoding the symbols in $J$ form a subset to which the AGHP space is $\varepsilon_{q,t}$-close to uniform; decoding preserves this accuracy. ∎
Remark.
The choice $\varepsilon_{q,t}:=(qt|C|)^{-c}$ is far smaller than needed in every later bound, yet keeps $|\mathcal{S}|$ polynomial in all input parameters. Hence we dispense with a user-supplied accuracy parameter for the rest of the paper.
2.2 The Relaxed Interlace Operation
Definition 10 (Relaxed Interlace Operation).
Given a base matrix $M:R\times C\to\{0,1\}$ and a balanced set $\mathcal{S}\subseteq C^{q}$ as in Definition 9, define the relaxed interlace of order $q$ by
$M\left\langle q,\mathcal{S}\right\rangle\;:\;([q]\times R)\times\mathcal{S}\;\longrightarrow\;\{0,1\},\qquad M\left\langle q,\mathcal{S}\right\rangle\left((i,r),\mathbf{c}\right)\;:=\;M\left(r,c_{i}\right).$
Only the $|\mathcal{S}|$ allowable column tuples are used, keeping the domain size
$|[q]\times R|\cdot|\mathcal{S}|\;=\;q\,|R|\,\cdot\,2^{\,O\!\left(\log\log(q\ell)+t\ell+\log(1/\varepsilon)\right)}$
polynomial in the natural parameters instead of exponential in $q$.
Discussion.
Intuitively, $M\left\langle q,\mathcal{S}\right\rangle$ behaves like the full interlace operation but draws its $q$-tuple of columns from a carefully‑balanced “reservoir” that is almost $t$-wise independent. Lemma 8 ensures that any protocol whose analysis touches at most $t$ of the $q$ interlaced positions at a time cannot distinguish the relaxed construction from the ideal, up to statistical error $\varepsilon$. Thus we inherit the complexity lower bounds and control over protocol behaviour in Section 3 while keeping the instance size polynomial in $p,q,t$ and $\log(1/\varepsilon)$.
2.3 Structural Lemmas on Relaxed Interlace
Lemma 9 (Balancing–by–Blocks).
Let $M$ be an $m\times n$ matrix and fix an integer $q\geq 1$. Write the row universe of the $q$-fold interlace $M\langle q\rangle$ as
$\widehat{R}\;=\;[q]\times R,\qquad m:=|R|.$
For a parameter $x\in(0,1]$ set $T:=\lceil mx\rceil$ and call a block $\{i\}\times R$ full if it contains at least $T$ rows of the subset under consideration.
Let $R^{\prime}\subseteq\widehat{R}$ be arbitrary and define the current row-density
$\beta\;:=\;\frac{|R^{\prime}|}{q\,m}\;\in[0,1].$
  1. (i)
    If $\beta<x$ then $R^{\prime}$ contains fewer than $qT$ rows, at least one of the $q$ blocks is not full; the only $T$-equipartitioned sub-selection is $R^{\prime\prime}=\varnothing$.
  2. (ii)
    Suppose $\beta\geq x$. Define $k:=\left\lceil q\,\frac{\beta-x}{1-x}\right\rceil$. Then there exists a block set $J\subseteq[q]$ of size $k$ and a subset
    $R^{\prime\prime}\;:=\;R^{\prime}\cap\left(J\times R\right)\;\subseteq\;R^{\prime}$
    such that
    $\forall i\in J:\quad\left|\,R^{\prime\prime}\cap(\{i\}\times R)\right|\;\geq\;T.$
    Thus $R^{\prime\prime}$ is $(J,T)$-equipartitioned, and $|J|$ satisfies the lower bound above.
Columns remain untouched; the lemma concerns rows only.
Proof.
For each block $i\in[q]$ write $r_{i}:=|R^{\prime}\cap(\{i\}\times R)|$ and let $Q^{\prime}:=\{\,i\in[q]\mid r_{i}<T\}$ be the set of deficient blocks.
Case $\beta<x$.
Then $|R^{\prime}|<qT$. Every block holds <T<T rows, so no non-empty $T$-equipartitioned subset exists; we must output $R^{\prime\prime}=\varnothing$.
Case $\beta\geq x$.
Because $|R^{\prime}|=q\beta m\geq qT$, at least one block is full. We now bound $|Q^{\prime}|$. Each deficient block is missing at least $T-r_{i}\geq m-T$ rows relative to the quota $T$, hence
$qT-|R^{\prime}|=\sum_{i\in[q]}(T-r_{i})\geq\sum_{i\in Q^{\prime}}(T-r_{i})\geq|Q^{\prime}|\,(m-T).$
Rearranging,
$|Q^{\prime}|\;\leq\;\frac{qT-|R^{\prime}|}{m-T}=q\,\frac{x-\beta}{1-x}.$
Set $J:=[q]\setminus Q^{\prime}$ so that
$|J|=q-|Q^{\prime}|\;\geq\;q-q\,\frac{x-\beta}{1-x}=q\,\frac{\beta-x}{1-x}\;\Longrightarrow\;|J|\geq k$
by the ceiling definition of $k$.
Constructing $R^{\prime\prime}$.
Pick any subset $\widehat{J}\subseteq J$ of exactly $k$ indices and let $R^{\prime\prime}:=R^{\prime}\cap(\widehat{J}\times R).$ For every $i\in\widehat{J}\subseteq J$ we have $r_{i}\geq T$, hence the quota condition $|R^{\prime\prime}\cap(\{i\}\times R)|\geq T$ is satisfied. No columns were altered, completing the proof. ∎
Lemma 10 (From relaxed to classical interlace).
Fix a base matrix $M:R\times C\to\{0,1\}$ with $|R|=m$ rows. Let $q\geq t\geq 1$ and let $\mathcal{S}=S_{q,t}(C)\subseteq C^{q}$ be the (q,t)(q,t)‑balanced column multiset of Definition 9 with accuracy $\varepsilon:=\varepsilon_{q,t}$. Denote by
$N\;:=\;M\left\langle q,\mathcal{S}\right\rangle\;:\;([q]\times R)\times\mathcal{S}\;\longrightarrow\;\{0,1\}.$
Fix a subset $Q\subseteq[q]$ of size $t$ and set $T:=\left\lceil m\,x\right\rceil$ for some parameter $x\in(0,1]$. Let
$R^{\prime}\subseteq[q]\times R\quad\text{and}\quad C^{\prime}\subseteq\mathcal{S}$
satisfy
  1. (a)
    $R^{\prime}$ is a (Q,T)(Q,T)‑equipartition in the sense of Definition 5, and
  2. (b)
    $\frac{|C^{\prime}|}{|\mathcal{S}|}\;>\;y$ for some $y\in(0,1]$.
Define
$y^{\prime}\;:=\;1-\left(1+\varepsilon\right)\,(1-y)\;=\;y-\varepsilon(1-y)\;=\;(1+\varepsilon)\,y-\varepsilon.$
Then there exist $R^{\ast}\subseteq R^{\prime}$ and $C^{\ast}\subseteq C^{\prime}$ such that
$\varepsilon\!\left(N;R^{\ast},C^{\ast}\right)\;\in\;M\left[t,x,\,y^{\prime}\right].$
In other words, the restriction of the relaxed interlace to $(R^{\prime},C^{\prime})$ contains, as a subgame, a member of the classical bracket family $M\left[t,x,y^{\prime}\right]$.
Proof.
Step 1: trimming the rows. Because $R^{\prime}$ is (Q,T)(Q,T)‑equipartitioned and $|Q|=t$, choose exactly $T$ rows from each block $\{q\}\times R$ with $q\in Q$ and discard every other row. Denote the surviving set by $R^{\ast}\subseteq R^{\prime}.$ Then $R^{\ast}$ is an exact (Q,T)(Q,T)‑equipartition, hence $\left|R^{\ast}\right|=tT$ and condition (i) of Definition 8 with parameter $x$ is met.
Step 2: projecting the columns. Let $\pi_{Q}:\mathcal{S}\to C^{t}$ be the natural projection that keeps the $t$ coordinates indexed by $Q$. Write $D:=\pi_{Q}(C^{\prime})\subseteq C^{t}.$ Because $C^{\prime}$ is a sub‑multiset of $\mathcal{S}$, $D$ inherits multiplicities. By the balance Lemma 8, the uniform distribution on $\mathcal{S}$ is $\varepsilon$‑close to tt‑wise independence; equivalently, every pattern in $C^{t}$ occurs in $\mathcal{S}$ with frequency $\frac{1}{|C|^{\,t}}\pm\varepsilon$. A standard averaging argument shows that at most an $\varepsilon$‑fraction of the tuples in $C^{t}$ are undersupplied compared with the ideal uniform frequency. Keeping only the well‑supplied patterns,
$D^{\ast}\;:=\;\left\{\,\mathbf{c}\in C^{t}\;\big{|}\;\text{frequency of $\mathbf{c}$ in $C^{\prime}$ }\geq\left(1-\varepsilon\right)\frac{|C^{\prime}|}{|C^{t}|}\right\},$
we lose at most an $\varepsilon$‑fraction of columns from $C^{\prime}$. More precisely,
$|D^{\ast}|\;\geq\;|C^{t}|\;\left(y-\varepsilon(1-y)\right)\;=\;|C^{t}|\,y^{\prime}.$
Step 3: lifting back to $\mathcal{S}$. For every pattern $\mathbf{c}\in D^{\ast}$ keep one representative column in $C^{\prime}$ whose projection is $\mathbf{c}$; call the resulting set $C^{\ast}\subseteq C^{\prime}$. By construction $|C^{\ast}|=|D^{\ast}|\geq|C^{t}|\,y^{\prime}$.
Step 4: identifying the subgame. Re‑index the rows in $R^{\ast}$ as $([t]\times R)\ni(i,r)$ with $i$ the coordinate in $Q$ (keeping the original order), and re‑index each column in $C^{\ast}$ by its QQ‑projection in $C^{t}$. Under that relabelling the submatrix $N^{\ast}:=\varepsilon\!\left(N;R^{\ast},C^{\ast}\right)$ is exactly the classical interlace $M\langle t\rangle$ restricted to $R^{\ast}_{\text{canon}}\times D^{\ast}.$ The row set is an exact TT‑equipartition, and $|D^{\ast}|=\lceil|C|^{t}\,y^{\prime}\rceil$ by the previous step, hence $N^{\ast}\in M\left[t,x,y^{\prime}\right].$
Therefore $N^{\ast}$ is a member of the target bracket family and, by construction, a subgame of the original restriction $\varepsilon(N;R^{\prime},C^{\prime})$, completing the proof. ∎
3 Protocol Control
In this section, we show how the (Relaxed Interlace Operation). preserves relevant properties of the (Interlacing Operation). when it comes to the behaviour of communication protocols.
In Mackenzie and Saffidine [4], lower bounds on the (Interlacing Operation). were derived as shown in the lemmas in section 1.3. We will first show that the relevant lower bounds can be preserved.
Definition 11 ($(\delta,b)$-robust matrix).
For an integer $b\geq 0$, $\delta\in(0,\frac{1}{\sqrt{2}}-\frac{1}{2}]$ and a Boolean matrix $M$, assume that $M$ satisfies the following robustness conditions
$D\left(M\right)$   $\;\;\geq\;\;1$ (Condition 1)
$D\left(M\left[1,2^{-b},\tfrac{1}{2}+\delta\right]\right)$ $\;\;\geq\;\;D\left(M\right)$ (Condition 2)
Then we call $M$ a $(\delta,b)$-robust matrix.
Lemma 11.
Given a $(\delta,b)$-robust matrix $M$ (definition 11), we have
$D\left(M\left[1,2^{-b},\tfrac{1}{8}+\tfrac{\delta}{4}\right]\right)$ $\;\;\geq\;\;D\left(M\right)-2\;\;\geq\;\;1$ (Condition 3)
$D\left(M\left[1,2^{-b},\tfrac{1}{4}+\tfrac{\delta}{2}\right]\right)$ $\;\;\geq\;\;D\left(M\right)-1$ (Condition 4)
$D\left(M\left[2,2^{-b+1},(\tfrac{1}{2}+\delta)^{2}\right]\right)$ $\;\;\geq\;\;D\left(M\right)+1.$ (Condition 5)
Proof.
Condition 3 By Lemma 6, we have
$D\left(M\left[1,2^{-b},\tfrac{1}{8}+\tfrac{\delta}{4}\right]\right)$
$\geq\;\;$ $D\left(M\left[1,2^{-b},\tfrac{1}{2}+\delta\right]\right)-2$
$\geq\;\;$ $1$
Condition 4: By Lemma 6, we have
$D\left(M\left[1,2^{-b},\tfrac{1}{4}+\tfrac{\delta}{2}\right]\right)$
$\geq\;\;$ $D\left(M\left[1,2^{-b},\tfrac{1}{2}+\delta\right]\right)-1$
$\geq\;\;$ $D\left(M\right)-1$
Condition 5: Since the communication complexity of a matrix is at least one greater than that of both submatrices obtained by a single partitioning, we consider the partition of $D\left(M\left[2,2^{-b+1},\tfrac{1}{4}+\tfrac{\delta}{2}\right]\right)$ along both rows and columns.
Case 1: Row Partition:
$D\left(M\left[2,2^{-b+1},(\tfrac{1}{2}+\delta)^{2}\right]\right)$
$\geq\;\;$ $D\left(M\left[1,2^{-b+1},\tfrac{1}{2}+\delta\right]\right)$ (By Lemma 2)
$\geq\;\;$ $D\left(M\left[1,2^{-b},\tfrac{1}{2}+\delta\right]\right)+1$ (Partition along rows)
$\geq\;\;$ $D\left(M\right)+1.$ (By Condition 2)
Case 2: Column Partition:
$D\left(M\left[2,2^{-b+1},(\tfrac{1}{2}+\delta)^{2}\right]\right)$
$\geq\;\;$ $D\left(M\left[2,2^{-b+1},\tfrac{(\tfrac{1}{2}+\delta)^{2}}{2}\right]\right)+1.$ (Partition along columns)
Then consider a further row or column partition:
Case 2a: Row Partition:
$D\left(M\left[2,2^{-b+1},\tfrac{(\tfrac{1}{2}+\delta)^{2}}{2}\right]\right)+1$
$\geq\;\;$ $D\left(M\left[1,2^{-b+1},\tfrac{\tfrac{1}{2}+\delta}{\sqrt{2}}\right]\right)+1$ (By Lemma 2)
$\geq\;\;$ $D\left(M\left[1,2^{-b},\tfrac{\tfrac{1}{2}+\delta}{\sqrt{2}}\right]\right)+2$ (Partition along rows)
$\geq\;\;$ $(D\left(M\right)-1)+2.$ (By Condition 4 and Condition 5)
Case 2b: Column Partition:
$D\left(M\left[2,2^{-b+1},\tfrac{(\tfrac{1}{2}+\delta)^{2}}{2}\right]\right)+1$
$\geq\;\;$ $D\left(M\left[2,2^{-b+1},\tfrac{(\tfrac{1}{2}+\delta)^{2}}{4}\right]\right)+2.$ (Partition along columns)
Again, we can partition the matrix along rows or columns:
Case 2b1: Row Partition:
$D\left(M\left[2,2^{-b+1},\tfrac{(\tfrac{1}{2}+\delta)^{2}}{4}\right]\right)+2$
$\geq\;\;$ $D\left(M\left[1,2^{-b+1},\tfrac{\tfrac{1}{2}+\delta}{2}\right]\right)+2$ (By Lemma 2)
$\geq\;\;$ $D\left(M\left[1,2^{-b},\tfrac{\tfrac{1}{2}+\delta}{2}\right]\right)+3$ (Partition along rows)
$\geq\;\;$ $(D\left(M\right)-1)+3.$ (By Condition 4)
Case 2b2: Column Partition:
$D\left(M\left[2,2^{-b+1},\tfrac{(\tfrac{1}{2}+\delta)^{2}}{4}\right]\right)+2$
$\geq\;\;$ $D\left(M\left[2,2^{-b+1},\tfrac{(\tfrac{1}{2}+\delta)^{2}}{8}\right]\right)+3$ (Partition along columns)
$\geq\;\;$ $D\left(M\left[1,2^{-b+1},\tfrac{\tfrac{1}{2}+\delta}{2\sqrt{2}}\right]\right)+3$ (By Lemma 2)
$\geq\;\;$ $(D\left(M\right)-2)+3.$ (By Condition 3)
3.1 Lower Bounding the Complexity
Lemma 12.
Let $M$ be a $m\times p$, $(\delta,b)$-robust matrix (definition 11) with $3\leq D\left(M\right)\leq b$, and let $p\geq 0$ be an integer.
Then
$D\left(M\left[2^{p}+1,2^{p-b+1},\tfrac{1}{2}+\delta\right]\right)\geq D\left(M\right)+p+1\geq D\left(M\left\langle 2^{p+1}\right\rangle\right).$
Proof.
Throughout this argument write
$y\;:=\;\tfrac{1}{2}+\delta,\qquad x_{i}\;:=\;2^{i-b}.$
We proceed by induction on $p$ to establish the lemma.
Base case $\boldsymbol{p=0}$. By robustness Condition 1,
$D\left(M\left[2,2^{-b+1},y\right]\right)$
$\geq\;\;$ $D\left(M\left[2,x_{1},y\right]\right)$ (By the definition of $x_{i}$)
$\geq\;\;$ $D\left(M\left[1,\tfrac{x_{1}}{2},y\right]\right)+1$ (By Lemma 3)
$\geq\;\;$ $D\left(M\left[2,2^{-b},y\right]\right)$ (By the definition of $x_{i}$)
$\geq\;\;$ $D\left(M\right)+1$ (By Condition 2)
$\geq\;\;$ $D\left(M\left\langle 2\right\rangle\right).$
Note the Lemma 3 is applied with $p=1$, $\delta=0$ and $x=\tfrac{x_{0}}{2}$ and $D\left(M\left[1,x_{0},y/4\right]\right)\geq 1$ holds due to Condition 3.
Inductive step. Assume the lemma holds for all $p<p^{\prime}$, and consider the case $p=p^{\prime}$.
$D\left(M\left[2^{p}+1,2^{p-b},y\right]\right)$
$\geq\;\;$ $D\left(M\left[2^{p}+1,x_{p},y\right]\right)$ (By the definition of $x_{i}$)
$\geq\;\;$ $D\left(M\left[2^{p-1}+1,x_{p-1},y\right]\right)+1$ (By Lemma 3)
$\geq\;\;$ $D\left(M\right)+p+1$ (By the inductive hypothesis)
$\geq\;\;$ $D\left(M\left\langle 2^{p+1}\right\rangle\right).$
Note the Lemma 3 is applied with $p=p$, $\delta=1$ and $x=x_{p-1}$ and $D\left(M\left[2^{p},x_{p},y/4\right]\right)\geq 1$ holds due to Lemma 1 and Condition 3. ∎
Lemma 13 (General Hard Seed Lemma).
Given a $(\delta,b)$-robust matrix $M$ with $D\left(M\right)\geq 3$, there exists $T\in\mathbb{N}$ such that for all $T\leq t\leq 2^{b}$ where $t$ is a power of $2$, and $1\leq\Delta\leq\log{t}$, we have
$D\left(M\left[\dfrac{(2^{\Delta-1}+2)t}{2^{\Delta}},2^{-b}t,2^{-\,2^{0.49\sqrt{\log t}}}\right]\right)\;\;\geq\;\;D\left(M\right)\;+\;\log t.$
Proof.
Start from the seed instance $M\left[2^{\Delta-1}+1,2^{-b+\Delta},\frac{1}{2}+\delta\right]$, whose complexity is at least $D\left(M\right)+\Delta$ by Lemma 12. We will apply Lemma 4 to this seed.
Parameter choice.
$p:=2^{\Delta-1}+1,\quad x:=2^{-b+\Delta},\quad y:=\frac{1}{2}+\delta,\quad k:=\log t-\Delta,\quad s:=\left\lceil\frac{\sqrt{\log t}}{\log\log t}\right\rceil,\quad\rho:=\sqrt{\log t}\;.$
Checklist for Lemma 4.
  1. (a)
    $s\leq k$. Since $k=\log t-\Delta$ and $s\in o(\log t)$, for sufficiently large $t$ we have $k\geq s$.
  2. (b)
    $0<x\leq 2^{-k}$. We have $x=2^{-b+\Delta}$ while $2^{-k}=2^{-\log t+\Delta}=\frac{2^{\Delta}}{t}$. From the condition that $t\leq 2^{b}$ we get that $2^{-k}=\frac{2^{\Delta}}{t}\geq 2^{-b+\Delta}=x$.
  3. (c)
    $0<y\leq 1$ is immediate.
  4. (d)
    $\rho>2$. Because $\rho=\sqrt{\log t}$ and $\sqrt{\log t}\to\infty$ with $t$, the inequality holds for all $t\geq T$.
  5. (e)
    $(\rho-1)^{k}\leq\rho^{\,k-s}$. Take the logarithms:
    $k\log(\rho-1)\;\leq\;(k-s)\log\rho\quad\Longleftrightarrow\quad s\log\rho\;\leq\;k\left[\log\rho-\log(\rho-1)\right].$
    Since $\log\rho-\log(\rho-1)=-\log\!\left(1-\tfrac{1}{\rho}\right)\leq\tfrac{1}{\rho}$, a sufficient condition is
    $s\;\leq\;\frac{k}{\rho\,\log\rho}\quad\Longrightarrow\quad k\;\geq\;s\,\rho\log\rho.$
    Now recall the parameter scales: $k=\log{t}-\Delta$ while $s=\left\lceil\frac{\sqrt{\log t}}{\log\log t}\right\rceil$ and $\rho=\sqrt{\log t}$. Hence the right‑hand side of ()(*) is
    $s\,\rho\log\rho$ $=\left\lceil\frac{\sqrt{\log t}}{\log\log t}\right\rceil\sqrt{\log t}\log{\sqrt{\log t}}$
    $\leq\left(\frac{\sqrt{\log t}}{\log\log t}+1\right)\sqrt{\log t}\frac{1}{2}\log{\log t}$
    $=\frac{1}{2}\log t+\frac{1}{2}\sqrt{\log t}\log\log t$
    $<\log t-\Delta=k.$ (for sufficiently large $t$)
    hence the desired inequality $(\rho-1)^{k}\leq\rho^{\,k-s}$ holds automatically for all sufficiently large $t$.
  6. (f)
    $D\left(M\left[p,x,y/4\right]\right)\geq 1$. Robustness directly gives $D\left(M\left[1,x,y/4\right]\right)\geq 1$ (Condition 3). Because complexity cannot decrease when we replace the single‑fold interlace by a $2^{\Delta-1}+1$‑fold one (apply Lemma 1 with $p^{\prime}=1,\;p=2^{\Delta-1}+1$), the required lower bound also holds for $p=2^{\Delta-1}+1$.
Thus every hypothesis of Lemma 4 is satisfied.
Applying Lemma 4. We first make a few observations about the values obtained.
For every $\rho>2$
$\left(\frac{\rho-1}{\rho-2}\right)^{s}\;=\;\left(1+\frac{1}{\rho-2}\right)^{s}\;\leq\;\exp\!\left(\frac{s}{\rho-2}\right),$
because $1+u\leq e^{u}$. Since $\frac{s}{\rho-2}\rightarrow 0$ as $t\rightarrow\infty$, we get
$\left(\frac{\rho-1}{\rho-2}\right)^{s}\;\leq\;\frac{2^{\Delta-1}+2}{2^{\Delta-1}+1}.$ (5)
For $y^{\rho^{s}}$, since $y=\frac{1}{2}+\delta\leq\frac{1}{\sqrt{2}}$ we observe that
$y^{\rho^{s}}$  $\;\leq\;2^{-\frac{1}{2}\,\rho^{s}}$
= $\;=\;2^{-\frac{1}{2}\left(\sqrt{\log t}\right)^{\left\lceil\frac{\sqrt{\log t}}{\log\log t}\right\rceil}}$
 $\;\leq\;2^{-\frac{1}{2}\left(\sqrt{\log t}\right)^{\frac{\sqrt{\log t}}{\log\log t}}}$
= $\;=\;2^{-\frac{1}{2}2^{\frac{\sqrt{\log t}}{\log\log t}\,\log\!\left(\sqrt{\log t}\right)}}$
= $\;=\;2^{-\frac{1}{2}2^{\frac{1}{2}\sqrt{\log t}\,\frac{\log\log t}{\log\log t}}}$
= $\;=\;2^{-\frac{1}{2}2^{\frac{1}{2}\sqrt{\log t}}}$
 $\;\leq\;2^{-\,2^{0.49\sqrt{\log t}}}\qquad(\text{for sufficiently large }t).$
Lemma 4 states that
$D\left(M\left[2^{k}\left(\frac{\rho-1}{\rho-2}\right)^{s}p,2^{k}x,y^{\rho^{s}}\right]\right)\geq k+D\left(M\left[p,x,y\right]\right).$
Substituting our parameters on the left hand side gives us
$D\left(M\left[2^{k}\left(\frac{\rho-1}{\rho-2}\right)^{s}p,2^{k}x,y^{\rho^{s}}\right]\right)$
$=\;$ $D\left(M\left[2^{\log t-\Delta}\left(\frac{\rho-1}{\rho-2}\right)^{s}(2^{\Delta-1}+1),2^{\log t-\Delta}2^{-b+\Delta},y^{\rho^{s}}\right]\right)$
$\leq\;$ $D\left(M\left[\frac{(2^{\Delta-1}+2)t}{2^{\Delta}},2^{-b}t,2^{-\,2^{0.49\sqrt{\log t}}}\right]\right),$
while the right hand side gives us
$k+D\left(M\left[p,x,y\right]\right)$
$=\;$ $\log t-\Delta+D\left(M\left[2^{\Delta-1}+1,2^{-b+\Delta},\frac{1}{2}+\delta\right]\right)$
$\geq\;$ $\log t+D\left(M\left[1,2^{-b},\frac{1}{2}+\delta\right]\right)$
$\geq\;$ $\log t+D\left(M\right),$
concluding the proof.
After presenting a generic way to obtain a provable ”hard seed” function from a generic $(\delta,b)$-robust function, we will now show that under the appropriate conditions one can extend any ”hard seed” lower bound with $t$ interlaced functions to show a lower bound on $rt$ interlaced functions interlaced via the relaxed interlace function.
Theorem 1 (Extension Theorem).
Given a $(\delta,b)$-robust matrix $M$, $r\geq 1$ be an integer, $t$ be a power of $2$ s.t. $t\leq 2^{b}$, $1\leq\Delta\leq\log{t}$, and given the following hard seed statement is true:
$D\left(M\left[\frac{(2^{\Delta-1}+2)t}{2^{\Delta}},2^{-b}t,h(t)\right]\right)\geq D\left(M\right)+\log{t}$ (6)
Where $h:\mathbb{N}\rightarrow(0,1]$. Writing $B:=D\left(M\right)+\log t+\left\lceil\log r\right\rceil$; $\varepsilon=\epsilon_{q,t}$ and assuming
$(1+\varepsilon)h^{\prime}(t)\cdot 2^{-B}r-\varepsilon$ $\geq h(t)$ (7)
$D\left(M\left[1,2^{-B},\left((1+\varepsilon)\,h^{\prime}(t)\,2^{-B}-\varepsilon\right)^{1/t}\right]\right)\;$  $\geq\;1,$ (8)
if we denote
$q:=rt,\quad q^{\prime}:=\frac{(2^{\Delta-1}+2)}{2^{\Delta}}rt,\quad\widehat{M}:=M\left\langle q,S_{q,t}(\mathsf{Cols}(M))\right\rangle,$
[…]
$R^{\prime}\text{ is }(Q,2^{-b}q|\mathsf{Rows}(M)|)\text{-equipartitioned with }|Q|=q^{\prime},$
$\frac{|C^{\prime}|}{|\mathsf{Cols}(\widehat{M})|}\geq h^{\prime}(t),$
we have that
$D\left(\varepsilon\left(\widehat{M};R^{\prime},C^{\prime}\right)\right)\geq D\left(M\right)+\log{t}+\left\lceil\log{r}\right\rceil.$
Proof.
Fix a deterministic protocol $\Pi$ for $\widetilde{M}:=\varepsilon\left(\widehat{M};R^{\prime},C^{\prime}\right)$ that communicates at most $B:=D\left(M\right)+\log t+\left\lceil\log r\right\rceil-1$ bits. We show $\Pi$ fails to reduce every leaf rectangle to complexity 0, which establishes the desired lower bound.
Dynamic invariant. Let $\mathsf{P}(i)$ be the statement: “After $i$ bits have been sent during the execution of $\Pi$ there exists a rectangle $R_{i}\times C_{i}$ in the current protocol partition for which $R_{i}$ is a $(Q,2^{-b})$-equipartition with $|Q|\geq t$.”
We may then distinguish two cases:
  1. a)
    $\Pi$ exhausts its budget $B$ without ever breaking the invariant.
  2. b)
    There exists $i\in[B]$ for which $P(i)$ is true and $P(i+1)$ is false.
Case a. Protocol spends its entire budget and the row–block invariant is never broken.
Let $\widetilde{M}=\varepsilon\left(\widehat{M};R_{i},C_{i}\right)$ be a rectangle that survives after the last communication bit.
  1. Step 1:
    Parameters of the surviving rectangle.
    $R_{i}\text{ is a }(Q,T_{0})\text{-equipartition},\qquad|Q|=t,\qquad T_{0}=\left\lceil|R|\,2^{-b}\right\rceil.$
    Because at most $B$ column bits were revealed,
    $\frac{|C_{i}|}{|\mathsf{Cols}(\widehat{M})|}\;\geq\;h^{\prime}(t)\,2^{-B}\;=:\;y.$
  2. Step 2:
[…]
    Apply the lemma with $t=|Q|$, $x=2^{-B}$, and the above $y$. It returns subsets $R^{\ast}\subseteq R_{i}$, $C^{\ast}\subseteq C_{i}$ and a subgame
    $N^{\ast}\;\in\;M\left[t,2^{-B},y^{\prime}\right],\qquad y^{\prime}\;=\;(1+\varepsilon)\,h^{\prime}(t)\,2^{-B}-\varepsilon.$
  3. Step 3:
    Projection to a single block.
    By Lemma 2,
    $D\left(N^{\ast}\right)\;\geq\;D\left(M\left[1,2^{-B},\left((1+\varepsilon)\,h^{\prime}(t)\,2^{-B}-\varepsilon\right)^{1/t}\right]\right).$
  4. Step 4:
    Seed assumption for one-block bracket.
    According to Equation 8,
    $D\left(M\left[1,2^{-B},\left((1+\varepsilon)\,h^{\prime}(t)\,2^{-B}-\varepsilon\right)^{1/t}\right]\right)\;\geq\;1.$
    Hence $D\left(N^{\ast}\right)\geq 1$.
Case b. The row–block invariant is broken at the first step i+1i{+}1 ($1\leq i<B$).
Let $\widetilde{M}_{i}=\varepsilon\left(\widehat{M};R_{i},C_{i}\right)$ be the unique rectangle that satisfies the invariant immediately before the (i+1)(i{+}1)-st bit is sent.
  1. Step 1:
    Row and column parameters of $\widetilde{M}_{i}$.
    By the invariant $R_{i}$ is a $(Q,T_{0})$-equipartition with
    $t=|Q|,\qquad T_{0}=\left\lceil|R|\,2^{-b}\right\rceil.$
    To break the invariant at step i+1i{+}1 the protocol must have asked at least $\left\lceil\log r\right\rceil$ row questions (each halves the number of viable blocks). Hence no more than $B-\left\lceil\log r\right\rceil$ bits have attacked the columns, so
    $\frac{|C_{i}|}{|\mathsf{Cols}(\widehat{M})|}\;\geq\;h^{\prime}(t)\;2^{-(B-\lceil\log r\rceil)}\;\geq\;h^{\prime}(t)\,r\,2^{-B}\;=:\;y.$
  2. Step 2:
[…]
    Apply the lemma to $\widetilde{M}_{i}$ with
    $t=|Q|,\qquad x=2^{-b},\qquad y=h^{\prime}(t)\,r\,2^{-B}.$
    It returns subsets $R^{\ast}\subseteq R_{i}$, $C^{\ast}\subseteq C_{i}$ and a subgame
    $N^{\ast}\;\in\;M\left[t,2^{-b},y^{\prime}\right],\qquad y^{\prime}=(1+\varepsilon)\,h^{\prime}(t)\,r\,2^{-B}-\varepsilon.$
    By condition (7) in the theorem we have $y^{\prime}\geq h(t)$.
  3. Step 3:
    Invoke the hard‑seed inequality.
    By the theorem’s hard‑seed hypothesis,
    $D\left(N^{\ast}\right)\;\geq\;D\left(M\right)+\log t.$
  4. Step 4:
    Compare with the bits still available.
    At time $i$ the protocol has already used $\geq\lceil\log r\rceil$ bits on rows, so at most
    $B-\lceil\log r\rceil\;=\;\left(D\left(M\right)+\log t+\lceil\log r\rceil-1\right)-\lceil\log r\rceil\;=\;D\left(M\right)+\log t-1$
    bits remain on any branch. Because
    $D\left(N^{\ast}\right)\;\geq\;D\left(M\right)+\log t\;>\;B-\lceil\log r\rceil,$
    the subtree below time $i$ cannot be fully resolved, which is a contradiction.
Therefore Case b is impossible, completing the proof of the Extension Theorem.
3.2 Near-Exact Separation of Interlaced Functions
In this section we will show that under the appropriate robustness conditions, an optimal protocol must separate interlaced functions nearly exactly. The requirements for this are slightly stronger than what was required for our complexity bounds, but Theorem 2 gives us the control on the protocol that we need to encode reductions.
Lemma 14.
Fix integers $b\geq 1$, $r\geq 1$ and let $t\leq 2^{b}$ be a power of two. Choose $0<\delta\leq\tfrac{1}{\sqrt{2}}-\tfrac{1}{2}.$
Assume $M$ is $(\delta,b)$‑robust and that
$D\left(M\left[2,2^{-\log{r}-\log{t}},\left((1+\varepsilon)\,2^{-3\log{r}-3\log{t}}-\varepsilon\right)^{1/t}\right]\right)\;$ $\geq\;D\left(M\right)+1,$ (9)
Let $q:=rt$ and write $\widehat{M}:=M\left\langle q,S_{q,t}(\mathsf{Cols}(M))\right\rangle.$ Pick $C^{\prime}$ with
$\dfrac{|C^{\prime}|}{|\mathsf{Cols}(\widehat{M})|}\geq\tfrac{1}{2}+\delta.$ (10)
Put $N:=\varepsilon\left(\widehat{M};\mathsf{Rows}(M),C^{\prime}\right)$ and $B^{\prime}:=\log t+\lceil\log r\rceil$.
For every integer $c\geq 1$, every deterministic protocol on $N$ that sends at most $B^{\prime}$ bits but spends fewer than $B^{\prime}-c$ of them on the outer block index leaves a transcript leaf rectangle $R_{c}\times C_{c}$ such that
$D\left(\varepsilon\left(\widehat{M};R_{c},C_{c}\right)\right)\;\geq\;D\left(M\right)+1.$
Hence any protocol attaining the tight bound $D\left(M\right)+\log t+\lceil\log r\rceil$ must devote all $B^{\prime}$ block bits to isolating the $q=rt$ interlaced games.
Proof.
Let $B^{\prime}=\log t+\lceil\log r\rceil$ and fix an integer $1\leq c\leq B^{\prime}$. Write $s:=B^{\prime}-c$ for the number of questions the protocol does ask about the outer block index.
Step 1: shape of the rectangle after $s$ block bits.
The analysis separates rows and columns.
Row side. Since $q=rt$ we can lower bound it by $2^{B^{\prime}-1}+1$. Since $B^{\prime}-c$ bits are spent separating the outer indices, at least one element of the partition must contain $\left\lceil\frac{2^{B^{\prime}-1}+1}{2^{B^{\prime}-c}}\right\rceil$ outer indices with at least $2^{-B^{\prime}}|\mathsf{Rows}(M)|$ inner indices present. Since $c\leq B^{\prime}$ this simplifies to $2^{c-1}+1$ outer indices that must be contained. This means that we can find a $Q,T$-equipartition $R_{c}$ with $|Q|\geq 2^{c-1}+1$ and $T\geq 2^{-B^{\prime}}$.
Column side. Each of the $c$ ”wasted bits” halve the current column set, so the surviving columns satisfy
$|C_{c}|\;\geq\;\frac{|C^{\prime}|}{2^{c}}\;\geq\;\frac{(\tfrac{1}{2}+\delta)\,|\mathsf{Cols}(\widehat{M})|}{2^{\,c}},$ (11)
where the last inequality uses condition (10).
Extraction. If the protocol does not separate all outer ranks, we obtain rectangle $R_{c}\times C_{c}$. Denoting $g(c,t):=\min{\left(2^{c}-1+1,t\right)}$, we may use Lemma 10 to infer
$N_{c}:=\varepsilon\left(\widehat{M};R_{c},C_{c}\right)\;\in\;M\left[g(c,t),\,2^{-B^{\prime}}\,,\displaystyle y_{c}:=\frac{\frac{1}{2}+\delta}{2^{\,c}}\right].$ (12)
Step 2: project onto two blocks (Condition 5).
Apply the (Extended Maximum Projection Lemma). with $l=2$:
$N_{c}^{\prime}\in M\left[2,2^{-\log{r}-\log{t}},\left(y_{c}\right)^{\,\frac{2}{{g(c,t)}}}\right].$
$\left(\dfrac{\frac{1}{2}+\delta}{2^{c}}\right)^{\frac{2}{g(c,t)}}$ $=\left(\frac{1}{2}+\delta\right)^{\frac{2}{g(c,t)}}\cdot 2^{-\frac{2c}{g(c,t)}}$
Given that $\frac{2}{g(c,t)}$ tends to zero the first term in the product goes to $1$ whilst
$2^{\frac{-2c}{g(c,t)}}\geq\left(2^{-2\log{r}-2\log{t}}\right)^{\frac{1}{t}}\geq\left((1+\varepsilon)\,2^{-3\log{r}-3\log{t}}-\varepsilon\right)^{\frac{1}{t}}.$
Therefore, condition 9 implies
$D\left(N_{c}^{\prime}\right)\;\geq\;D\left(M\right)+1.$
Step 3: monotonicity back to the protocol leaf. Because $N_{c}^{\prime}$ is a submatrix of the leaf rectangle $\widetilde{R}\times\widetilde{C}$, the Monotonicity Lemma 1 gives
$D\left(\varepsilon\left(\widehat{M};R_{c},C_{c}\right)\right)\;\geq\;D\left(N_{c}^{\prime}\right)\;\geq\;D\left(M\right)+1.$
This gives our gap. ∎
Theorem 2 (Near-Exact Separation Theorem).
Given a $(\delta,b)$-robust matrix $M$ such that
$D\left(M\left[2,2^{-\log{r}-\log{t}},\left((1+\varepsilon)\,2^{-\log{r}-\log{t}}-\varepsilon\right)^{1/t}\right]\right)\;$ $\geq\;D\left(M\right)+1,$ (13)
[…]
$N\in\varepsilon\left(\widehat{M};R,C\right).$
Then after $rt$ bits the elements of the partition must all be of the form $R^{\prime}\times\mathsf{Cols}(N)$ where $R^{\prime}\subseteq\mathsf{Rows}(N)$ is a $(Q,|\mathsf{Rows}(M)|(1-2^{-b}rt))$-equipartition for some $Q\subseteq[rt]$ and $|Q|=1$.
Proof.
Using Lemma 14, we may assume that the number of ”wasted bits” $c$ is 0, and hence no bit is spent partitioning the columns. From Condition 5, it we cannot have an element with $2^{-b}|\mathsf{Rows}(N)|$ rows from $2$ different outer indices. Since the number of outer indices equals the number of protocol leaves after rtr\,t bits, for any given outer index $i$ there is exactly one leaf $L_{i}=R_{i}\times C$ containing more than $|\mathsf{Rows}(N)|2^{-b}$ of the form $(i,e)$ with $e\in Rows(M)$. Since all rows must be mapped to a leaf, and there are rtr\,t leaves, this implies that for all $i$, $L_{i}=R_{i}\times C$ where $R_{i}$ contains at least $|\mathsf{Rows}(M)|\left(1-2^{-b}rt\right)$ rows of the form $(i,e)$ with $e\in\mathsf{Rows}(M)$. ∎
4 Reduction from $\{0,1\}$-Vector Bin Packing
In this section, we describe a reduction from the $d$-Dimension Vector Bin Packing problem to the Communication Complexity problem by leveraging the properties of the (Interlacing Operation). and the (Relaxed Interlace Operation)..
The $\{0,1\}$-$d$-Dimension Vector Bin Packing problem is defined as follows:
$\{0,1\}$-$d$-Dimension Vector Bin Packing
Input: A set of $n$ vectors $\{v_{1},\ldots,v_{n}\}$ from $\{0,1\}^{d}$ and $m$ the number of bins
Question: Is there a partition of the vectors into $m$ bins such for each bin $B_{i}$, $\lVert\sum_{v\in B_{i}}v\rVert_{\infty}\leq 1$?
The 0,1{0,1}-$d$-Dimension Vector Bin Packing problem is known to be NP-hard even when the number of bins $m=4$, via a reduction from the 4-Coloring problem. In this section, we will reduce $\{0,1\}$ $d$-Dimension Vector Bin Packing with $m=4$ to Communication Complexity.
4.1 Preprocessing
We begin with an instance of $n^{\prime}$ vectors $v^{\prime}_{1},\ldots,v^{\prime}_{n^{\prime}}\in\{0,1\}^{d^{\prime}}$ and $m=4$ bins; denote the $j$-th coordinate of $v^{\prime}_{i}$ by $v^{\prime}_{i,j}$. Without loss of generality, assume that for every coordinate $j\in[d^{\prime}]$ at most four vectors satisfy $v^{\prime}_{i,j}=1$; otherwise, by the pigeonhole principle, some bin would contain two ones in coordinate $j$, violating unit capacity and yielding an immediate NO instance.
Next, to handle boundary cases uniformly, we replicate each coordinate block threefold. Set $d:=3d^{\prime}$ and, for each $i\in[n^{\prime}]$, create three vectors $v^{(1)}_{i},v^{(2)}_{i},v^{(3)}_{i}\in\{0,1\}^{d}$ by
$v^{(1)}_{i}=(v^{\prime}_{i},\,0^{d^{\prime}},\,0^{d^{\prime}}),\qquad v^{(2)}_{i}=(0^{d^{\prime}},\,v^{\prime}_{i},\,0^{d^{\prime}}),\qquad v^{(3)}_{i}=(0^{d^{\prime}},\,0^{d^{\prime}},\,v^{\prime}_{i}),$
equivalently, for all $j\in[d^{\prime}]$,
$v^{(1)}_{i,j}=v^{\prime}_{i,j},\quad v^{(2)}_{i,d^{\prime}+j}=v^{\prime}_{i,j},\quad v^{(3)}_{i,2d^{\prime}+j}=v^{\prime}_{i,j},$
with all other coordinates zero. Let $n:=3n^{\prime}$ and index the new family as $\{\,v^{(b)}_{i}:i\in[n^{\prime}],\,b\in[3]\,\}$. Thus the transformed instance has $n=3n^{\prime}$ vectors in dimension $d=3d^{\prime}$, a structure we will exploit in the construction that follows.
It is easy to see this transformation preserves the original problem’s satisfiability. Let $l_{i,j}$ denote the $j$-th vector whose $i$-th element is 1, we may fill up 0s in each vector to make $d$ be an integer power of 2.
4.2 Construction
Definition 12 (Transpose).
The transpose of an $m\times n$ Boolean matrix $M$ is the matrix $M^{\top}\in\{0,1\}^{n\times m}$ defined by $M^{\top}(i,j)=M(j,i)$ for all $i\in[n]$ and $j\in[m]$.
High-level overview.
To encode a $\{0,1\}$ $d$-Dimension Vector Bin Packing instance as a two-party communication matrix, we iteratively build five Boolean matrices
$\phi,\;\mathcal{M}_{1},\;\mathcal{M}_{2},\;\mathcal{M}_{3},\;\text{and }\mathcal{M}_{4},$
each obtained from the previous one via interlace operations. At each stage we weave additional instance structure (vectors, dimensions, and bins) into the matrix, culminating in $\mathcal{M}_{4}$, whose rows and columns encode the target communication problem. We now describe the construction step by step; this will set up the lower-bound analysis that follows.
Stage 0: the seed matrix $\phi$.
We start from the $1\times 2$ seed
$\phi\;:=\;\begin{bmatrix}1&0\end{bmatrix}$,
which serves as the atomic pattern used throughout the interlacing.
Stage 1: first interlace $\mathcal{M}_{1}$.
Fix an integer $a>0$ and let $q_{1}=2^{a}-2$. Define the $(q_{1},t_{1})$-fold relaxed interlace of $\phi$ by
$\mathcal{M}_{1}\;:=\;\phi\left\langle q_{1},\,S_{q_{1},t_{1}}(\mathsf{Cols}(\phi))\right\rangle.$
Intuitively, $\mathcal{M}_{1}$ repeats the seed $q_{1}$ times, preparing a mechanism that later limits how many vectors can occupy the same position within a bin. We write $R_{1}:=\mathsf{Rows}(\mathcal{M}_{1})=[q_{1}]\times[1]$ and $C_{1}:=\mathsf{Cols}(\mathcal{M}_{1})$.
Stage 2: second interlace $\mathcal{M}_{2}$.
Next, we interlace the transpose of $\mathcal{M}_{1}$ with order $q_{2}$:
$\mathcal{M}_{2}\;:=\;\mathcal{M}_{1}^{\top}\left\langle q_{2},\,S_{q_{2},t_{2}}(R_{1})\right\rangle.$
This operation juxtaposes $q_{2}$ copies of $\mathcal{M}_{1}^{\top}$ so that each column picks a $q_{2}$-tuple of row indices from $R_{1}$, thereby distributing the instance’s dimensions across the construction. Set $R_{2}:=\mathsf{Rows}(\mathcal{M}_{2})=[q_{2}]\times C_{1}$ and $C_{2}:=\mathsf{Cols}(\mathcal{M}_{2})$.
Stage 3: third interlace $\mathcal{M}_{3}$.
We then apply a $4$-fold interlace to $\mathcal{M}_{2}^{\top}$:
$\mathcal{M}_{3}\;:=\;\mathcal{M}_{2}^{\top}\left\langle 4\right\rangle.$
Thus each column of $\mathcal{M}_{3}$ aggregates four candidates from $\mathcal{M}_{2}$, mirroring the requirement that vectors must be partitioned into $4$ bins. We denote $R_{3}:=\mathsf{Rows}(\mathcal{M}_{3})=[4]\times C_{2}$ and $C_{3}:=\mathsf{Cols}(\mathcal{M}_{3})=R_{2}^{\,4}$. This stage provides the scaffold onto which the concrete instance data will be attached.
Stage 4: instance-aware concatenation $\mathcal{M}_{4}$.
Finally, we augment $\mathcal{M}_{3}$ with rows that encode the input vectors. Formally, $\mathcal{M}_{4}$ keeps the first $4\lvert S_{q_{2},t_{2}}(\mathcal{M}_{1}^{\top})\rvert$ rows identical to those of $\mathcal{M}_{3}$, while duplicating each column of $\mathcal{M}_{3}$ $2^{5}=32$ times to facilitate a compact vector encoding. The remaining $n$ rows are in bijection with the instance vectors $\{v_{1},\dots,v_{n}\}$.
Let $R_{4}:=R_{3}\cup[n]$ and $C_{4}:=[32]\times R_{2}^{\,4}$. For $i\in R_{3}$ and $j=(j^{\prime},(j_{1},j_{2},j_{3},j_{4}))\in C_{4}$ (with $j^{\prime}\in[32]$ and $j_{1},j_{2},j_{3},j_{4}\in R_{2}$), we preserve $\mathcal{M}_{3}$:
$\mathcal{M}_{4}\!\left(i,\,(j^{\prime},(j_{1},j_{2},j_{3},j_{4}))\right)\;:=\;\mathcal{M}_{3}\!\left(i,\,(j_{1},j_{2},j_{3},j_{4})\right).$
For a vector row $i\in[n]$, we write
$\mathcal{M}_{4}\!\left(i,(j^{\prime},(j_{1},j_{2},j_{3},j_{4}))\right):=\begin{cases}\phi\left\langle 5\right\rangle(p,\,j^{\prime})&\text{if }\exists!\,p\in[4]:\;j_{1}=j_{2}=j_{3}=j_{4}\land j_{1}=l_{j_{1},p},\\ \phi\left\langle 5\right\rangle(5,\,j^{\prime})&\text{otherwise}.\end{cases}$
In words, when a column selects the same dimension in all four slots and points to the unique index $p\in[4]$ of the vector active at that dimension, we encode $p$ through a 5-fold interlace of the seed; otherwise we write a default symbol (the “5th” row of the same interlace). This ensures that valid bin assignments are reflected as structured patterns inside $\mathcal{M}_{4}$.
Remark.
The interlace operator is additive in the number of layers: $M\left\langle x\right\rangle\circ M\left\langle y\right\rangle=M\left\langle x+y\right\rangle$. Hence, fixing a column of $\mathcal{M}_{4}$ with $j_{1}=j_{2}=j_{3}=j_{4}=\hat{j}$ collapses the corresponding submatrix to an interlace of $\mathcal{M}_{1}$ with $\phi\left\langle 5\right\rangle$, i.e.,
$\phi\left\langle(2^{a}-2)+5\right\rangle\;=\;\phi\left\langle 2^{a}+3\right\rangle.$
Moreover, when exactly one of the four vectors is present for that dimension, the restriction becomes a submatrix of $\phi\left\langle 2^{a}\right\rangle$, yielding the one-bit gap we leverage later.
4.3 Correctness Proof
Table 3 gives a table of the parameters used to instantiate the reduction.
Parameter Value chosen in the paper Brief description / where first introduced
$c_{i}$ $i$ Fixed positive constants
[…]
$t_{1}$ $\left\lceil\,c_{64}\log d\,\right\rceil_{\!2}$ # independent blocks in Lemma 15
$q_{2}$ $\left\lceil d\right\rceil_{\!2}$ Outer order of the second relaxed interlace (Stage 2)
$t_{2}$ $\left\lceil\frac{c_{3}\log d}{\log\log d}\right\rceil_{\!2}$ # independent blocks used in Lemma 17
$q_{3}$ 44 Outer order of the third (classical) interlace (Stage 3)
$t_{3}$ 44 # blocks in the final 4‑fold interlace of $\mathcal{M}_{2}^{\top}$
$\delta$ $0.1$
$\Delta$ 55
$b_{0}$ $c_{64}\,\log\!\left(c_{64}\log d\right)$
$b^{\prime}_{1}$ $c_{3}\log d$
$b_{1}$ $c_{2}\log d$
$b^{\prime}_{2}$ $c_{4}\log\log d$
$b_{2}$ $c_{3}\log\log d$
$x$ $2^{-5}$ Balancing parameter in Lemma 9
Table 3: Key parameters, their explicit instantiation, and the first place they are used. All logarithms are base 2 unless stated otherwise.
Fixed integer constants.
For every integer $i\geq 1$ we set
$c_{i}\;:=\;i.$
Thus $c_{2}=2,\;c_{3}=3,$ and so on. We use these symbols to keep formulae compact when a literal integer would clutter the line.
Lemma 15.
Fix
$c_{64}=64,\qquad c_{1}=1,\qquad\Delta=5.$
Define
$t_{1}:=\left\lceil c_{64}\log d\right\rceil_{\!2},\quad b_{0}:=c_{64}\,\log\!\left(c_{64}\,\log d\right),\quad b^{\prime}_{1}:=c_{1}\,\log d=\log d.$
Then, with $p:=\dfrac{(2^{\Delta-1}+1)t_{1}}{2^{\Delta}},\;x:=2^{-b_{0}},\;y:=2^{-b^{\prime}_{1}},$ we have
$D\left(\phi\left[p,x,y\right]\right)\;\;\geq\;\;D\left(\phi\left\langle t_{1}\right\rangle\right).$
Proof.
Step 0. Numeric bounds. From the definition of $t_{1}$
$c_{64}\log d\;\leq\;t_{1}\;\leq\;2c_{64}\log d\quad\Longrightarrow\quad 64\log d\;\leq\;t_{1}\;\leq\;128\log d.$
With $\Delta=5$ we have
$p=\frac{17}{32}\,t_{1}\;\geq\;\frac{17}{32}\,(64\log d)=34\log d.$
Also $\log y=-b^{\prime}_{1}=-\log d$.
Step 1. Lower–bound the bracket using Lemma 5. Lemma 5 says
$D\left(\phi\left[p,x,y\right]\right)\;\geq\;\left\lceil\log\!\left(p+\log y\right)\right\rceil+1.$
Here $p+\log y\geq 34\log d-\log d=33\log d.$ Because $t_{1}/2\leq 64\log d$,
$p+\log y\;\geq\;\frac{t_{1}}{2}.$
Thus
$\left\lceil\log\!\left(p+\log y\right)\right\rceil\;\geq\;\left\lceil\log(t_{1}/2)\right\rceil\;=\;\left\lceil\log t_{1}\right\rceil-1,$
giving
$D\left(\phi\left[p,x,y\right]\right)\;\geq\;\left\lceil\log t_{1}\right\rceil.$ (14)
Step 2. Trivial upper bound for the plain interlace. In $\phi\left\langle t_{1}\right\rangle$ Bob can send the row index $i\in[t_{1}]$ using $\lceil\log t_{1}\rceil$ bits, after which both parties know the value of $\phi$. Hence
$D\left(\phi\left\langle t_{1}\right\rangle\right)\;\leq\;\left\lceil\log t_{1}\right\rceil.$ (15)
Step 3. Comparison. Combining (14) and (15) gives
$D\left(\phi\left[p,x,y\right]\right)\;\geq\;\left\lceil\log t_{1}\right\rceil\;\geq\;D\left(\phi\left\langle t_{1}\right\rangle\right),$
completing the proof. ∎
Lemma 16.
$\mathcal{M}_{1}^{\top}$ is $(\delta,b_{1})$-robust.
Proof.
Denote $p:=\dfrac{(2^{\Delta-1}+1)q_{1}}{2^{\Delta}}$. We first start by showing a lower bound on
$\varepsilon\left(\phi\left\langle q_{1}\right\rangle;R^{\prime}_{1},C^{\prime}_{1}\right)$
with $R^{\prime}_{1}\subseteq R_{1}$ and $C^{\prime}_{1}\subseteq C_{1}$ satisfying conditions
$R^{\prime}_{1}\text{ is a $(Q,2^{-b_{0}}|C_{1}|)$-equipartition with }|Q|=p$
$\frac{|C^{\prime}_{1}|}{|C_{1}|}\leq 2^{-b_{1}}.$
Proving a hard seed.
Lemma 15 gives us our hard seed
$D\left(\phi\left[p,2^{-b_{0}},2^{-b^{\prime}_{1}}\right]\right)\;\geq\;D\left(\phi\left\langle t_{1}\right\rangle\right).$
Extending the hard seed.
We feed the following parameters into Theorem 1
$M:=\phi,\quad t=t_{1},\quad h(t_{1})=2^{-c_{3}\log{d}},\quad\varepsilon=\varepsilon_{q_{1},t_{1}},$
along with our hard seed, which sets budget $B_{1}=\log{q_{1}}+1$. We can check that the two main conditions (7, 8) of Theorem 1 are satisfied.
From this we obtain that
$D\left(\varepsilon\left(\phi\left\langle q_{1},S_{q_{1},t_{1}}(\mathsf{Cols}(\phi))\right\rangle;R^{\prime}_{1},C^{\prime}_{1}\right)\right)\geq\log q_{1}+1=D\left(\mathcal{M}_{1}\right),$
showing our lower bound on $\varepsilon\left(\phi\left\langle q_{1}\right\rangle;R^{\prime}_{1},C^{\prime}_{1}\right)$.
Relating to the containing matrix.
For any matrix
$N\in\mathcal{M}_{1}\left[1,\frac{1}{2}+\delta,2^{-b_{1}}\right],$
there exists $\widetilde{R_{1}}$ and $C^{\prime}_{1}$ such that $\frac{|\widetilde{R_{1}}|}{|R_{1}|}\geq\frac{1}{2}+\delta$ and $\frac{|C^{\prime}_{1}|}{|C_{1}|}\geq 2^{-b_{1}}$. We may apply Lemma 9 (the balancing lemma) with balancing parameter $x=2^{-5}$ to show that from $\widetilde{R_{1}}$ we can obtain a $(Q,2^{-5}|C_{1}|)$-equipartition $R^{\prime}_{1}\subseteq\widetilde{R_{1}}$ with $|Q|=\frac{2^{\Delta-1}+1}{2^{\Delta}}q_{1}$.
Chaining it all together gives us
$D\left(\mathcal{M}_{1}^{\top}\left[1,\frac{1}{2}+\delta,2^{-b_{1}}\right]\right)$
$=\;$ $D\left(\mathcal{M}_{1}\left[1,2^{-b_{1}},\frac{1}{2}+\delta\right]\right)$
$\geq\;$ $D\left(\mathcal{M}_{1}\right),$
proving $(\delta,b_{1})$-robustness. ∎
Lemma 17.
$\mathcal{M}_{2}^{\top}$ is $(\delta,b_{2})$-robust.
Proof.
Denote $p:=\dfrac{(2^{\Delta-1}+1)q_{2}}{2^{\Delta}}$. We first start by showing a lower bound on
$\varepsilon\left(\mathcal{M}_{1}^{\top}\left\langle q_{2}\right\rangle;R^{\prime}_{2},C^{\prime}_{2}\right)$
with $R^{\prime}_{2}\subseteq R_{2}$ and $C^{\prime}_{2}\subseteq C_{2}$ satisfying conditions
$R^{\prime}_{2}\text{ is a $(Q,2^{-b_{1}}|C_{1}|)$-equipartition with }|Q|=p$
$\frac{|C^{\prime}_{2}|}{|C_{2}|}\leq 2^{-b_{2}}.$
Proving a hard seed.
We feed the following parameters into Lemma 13:
$M:=\mathcal{M}_{1}^{\top},\quad t:=t_{2},$
to get the statement that
$D\left(\mathcal{M}_{1}^{\top}\left[t_{2},2^{b_{1}+t_{2}},2^{0.49\sqrt{\log t_{2}}}\right]\right)\geq D\left(\mathcal{M}_{1}^{\top}\right).$
Extending the hard seed.
We feed the following parameters into Theorem 1
$M:=\mathcal{M}_{1}^{\top},\quad t=t_{2},\quad h(t_{2})=\left\lceil\frac{c_{3}\log d}{\log\log d}\right\rceil_{\!2},\quad\varepsilon=\varepsilon_{q_{2},t_{2}}.$
along with our hard seed, which sets budget $B_{2}=\log{q_{2}}+\log{q_{1}}+1$. We can check that the two main conditions (7, 8) of Theorem 1 are satisfied.
From this we obtain that
$D\left(\varepsilon\left(\phi\left\langle q_{2},S_{q_{2},t_{2}}(R_{1})\right\rangle;R^{\prime}_{2},C^{\prime}_{2}\right)\right)\geq\log q_{2}+\log q_{1}+1=D\left(\mathcal{M}_{2}\right).$
showing our lower bound on $\varepsilon\left(\phi\left\langle q_{2}\right\rangle;R^{\prime}_{2},C^{\prime}_{2}\right)$.
For any matrix
$N\in\mathcal{M}_{1}\left[1,\frac{1}{2}+\delta,2^{-b_{1}}\right],$
there exists $\widetilde{R_{1}}$ and $C^{\prime}_{1}$ such that $\frac{|\widetilde{R_{1}}|}{|R_{1}|}\geq\frac{1}{2}+\delta$ and $\frac{|C^{\prime}_{1}|}{|C_{1}|}\geq 2^{-b_{1}}$.
Relating to the containing matrix.
For any matrix
$N\in\mathcal{M}_{2}\left[1,\frac{1}{2}+\delta,2^{-b_{2}}\right],$
there exists $\widetilde{R_{2}}$ and $C^{\prime}_{2}$ such that $\frac{|\widetilde{R_{2}}|}{|R_{2}|}\geq\frac{1}{2}+\delta$ and $\frac{|C^{\prime}_{2}|}{|C_{2}|}\geq 2^{-b_{2}}$. We may apply Lemma 9 (the balancing lemma) with balancing parameter $x=2^{-5}$ to show that from $\widetilde{R_{2}}$ we can obtain a $(Q,2^{-5}|C_{1}|)$-equipartition $R^{\prime}_{2}\subseteq\widetilde{R_{2}}$ with $|Q|=\frac{2^{\Delta-1}+1}{2^{\Delta}}q_{2}$.
Chaining it all together gives us
$D\left(\mathcal{M}_{1}^{\top}\left[1,\frac{1}{2}+\delta,2^{-b_{2}}\right]\right)$
$=\;$ $D\left(\mathcal{M}_{1}\left[1,2^{-b_{2}},\frac{1}{2}+\delta\right]\right)$
$\geq\;$ $D\left(\mathcal{M}_{1}\right),$
proving $(\delta,b_{2})$-robustness.
Lemma 18.
Any matrix $N$ in $\mathcal{M}_{1}^{\top}\left[1,\left\lceil q_{2}\right\rceil_{\!2},\left((1+\varepsilon_{q_{2},t_{2}})\,h^{\prime}(t_{2})\,\left\lceil q_{2}\right\rceil_{\!2}-\varepsilon_{q_{2},t_{2}}\right)^{1/t_{2}}\right]$ is $(\delta,2)$-robust.
Proof Sketch.
Plugging in the values and using Lemma 13 gives the desired result. ∎
Lemma 19.
Any matrix $N$ in $\mathcal{M}_{2}^{\top}\left[1,\left\lceil q_{3}\right\rceil_{\!2},\left((1+\varepsilon_{q_{3},t_{3}})\,h^{\prime}(t_{3})\,\left\lceil q_{3}\right\rceil_{\!2}-\varepsilon_{q_{3},t_{3}}\right)^{1/t_{3}}\right]$ is $(\delta,2)$-robust.
Proof Sketch.
Plugging in the values and using Lemma 13 gives the desired result. ∎
Lemma 20 (Near–exact separation for $\mathcal{M}_{3}$).
Let $q_{3}=4$ and recall that the rows of $\mathcal{M}_{3}$ are indexed by $q_{3}]\times C_{2}$, where the outer block index is the first coordinate in $q_{3}]$. Fix the robustness parameter $b_{2}$ from Lemma 17.
Consider any deterministic protocol $\Pi$ for $\mathcal{M}_{3}$ and any branch of $\Pi$ after the first $\log q_{3}=2$ row questions (these are the only questions that can restrict the outer block). Let $R\times C$ be the current rectangle on that branch.
Then there exists a (unique) block $i^{\star}\in[q_{3}]$ such that
$\left|\,R\cap(\{i^{\star}\}\times C_{2})\,\right|\;\;\geq\;\;(1-2^{-b_{2}})\,|C_{2}|\quad\text{and}\quad\forall i\neq i^{\star}:\;\left|\,R\cap(\{i\}\times C_{2})\,\right|\;\;\leq\;\;2^{-b_{2}}\,|C_{2}|.$
In words, every rectangle surviving the first $\log q_{3}$ block bits is already “essentially single-block”: it contains (up to a $2^{-b_{2}}$ fraction) all rows of exactly one of the four interlaced copies of $\mathcal{M}_{2}$ and (up to the same tiny fraction) no rows from any other copy. Consequently, any protocol that hopes to be optimal must spend its entire $\log q_{3}$ row budget isolating the outer index of $\mathcal{M}_{3}$.
Proof.
Lemma 18 gives us the main condition (9) needed to apply Theorem 2 with
$M\;=\;\mathcal{M}_{1}^{\top},\qquad t\;=\;q_{3},\qquad r\;=\;1.$
Verifying every hypothesis.
  1. (i)
    By Lemma 17, $\mathcal{M}_{2}^{\top}$ is $(\delta,b_{2})$‑robust, so the “base‑matrix robustness” requirement of the theorem is met.
  2. (ii)
    $q_{3}$ is a power of two, and we fixed it so that $q_{3}\leq 2^{b_{2}}$, hence $t\leq 2^{b_{2}}$ as required.
  3. (iii)
    Condition (9) for $M=\mathcal{M}_{2}^{\top}$, $t=q_{3}$ and $r=1$ is exactly Lemma 18.
Invoke the theorem. Since $r=1$, we have
$B^{\prime}\;=\;\log t+\lceil\log r\rceil\;=\;\log q_{3}.$
Theorem 2 now tells us that after these $\log q_{3}$ row questions every surviving rectangle $R\times C\subseteq\mathcal{M}_{2}$ contains
  1. a)
    at least $(1-2^{-b_{2}}q_{2})\,|C_{2}|$ rows from a single outer block $\{i^{\star}\}\times C_{2}$, and
  2. b)
    at most $2^{-b_{2}}q_{2}\,|C_{2}|$ rows from any other block.
That is exactly the “near‑exact separation” property stated in Lemma 21. ∎
Lemma 21 (Near–exact separation for $\mathcal{M}_{2}$).
Let $q_{2}$ be the outer interlace order in the definition of $\mathcal{M}_{2}$ (rows are indexed by $q_{2}]\times C_{1}$) and retain the robustness parameter $b_{1}$ from Lemma 16. Let $\Pi$ be a deterministic protocol for $\mathcal{M}_{2}$.
After the first $\log q_{2}$ row questions of $\Pi$, each leaf-rectangle $R\times C$ in the induced partition satisfies
$\exists i^{\star}\in[q_{2}]:\;\left|\,R\cap(\{i^{\star}\}\times C_{1})\,\right|\;\;\geq\;\;(1-2^{-b_{1}}q_{2})\,|C_{1}|\quad\text{and}\quad\forall i\neq i^{\star}:\;\left|\,R\cap(\{i\}\times C_{1})\,\right|\;\;\leq\;\;2^{-b_{1}}q_{2}\,|C_{1}|.$
Thus every rectangle that survives the $\log q_{2}$ outer-block bits is already forced to (almost) one of the $q_{2}$ aggregated copies of $\mathcal{M}_{1}$. Any protocol attaining the tight $D\left(\mathcal{M}_{1}\right)+\log q_{2}$ bound must therefore devote all of its $\log q_{2}$ row bits to separating the $q_{2}$ blocks.
Proof.
Lemma 19 gives us the main condition (9) needed to apply Theorem 2. The proof follows the same arguments as 20 with the parameters one level down. ∎
Lemma 22 (Completeness of the embedded $\mathcal{M}_{1}$ copies in $\mathcal{M}_{2}$).
Let $q_{2}$ and $t_{2}$ be the parameters in the definition of
$\mathcal{M}_{2}\;=\;\mathcal{M}_{1}^{\top}\left\langle q_{2},\,S_{q_{2},t_{2}}(C_{1})\right\rangle,\qquad R_{2}=[q_{2}]\times C_{1},\quad C_{2}:=S_{q_{2},t_{2}}(C_{1}),$
and set $\varepsilon:=\varepsilon_{q_{2},t_{2}}$.
For every outer index $i\in[q_{2}]$ denote the $i$-th copy of $\mathcal{M}_{1}$ inside $\mathcal{M}_{2}$ by
$\mathcal{M}_{1}^{(i)}\;:=\;\mathcal{M}_{2}\!\left[\,\{i\}\times C_{1},\,C_{2}\right],\qquad R_{i}:=\{i\}\times C_{1}.$
(Multiplicity lower bound). By the almost–kk–wise independence of the reservoir $S_{q_{2},t_{2}}(C_{1})$ we have, for every $c\in C_{1}$,
$\left\{\,\mathbf{d}\in C_{2}:d_{i}=c\right\}\;\;\geq\;\;\left(\tfrac{1}{|C_{1}|}-\varepsilon\right)\,|C_{2}|\;=:\;\mu.$
(Fullness condition). Let $C^{\prime}\subseteq C_{2}$ be any column subset such that
$|C^{\prime}|\;\;\geq\;\;|C_{2}|-\mu\;=\;\left(1-\tfrac{1}{|C_{1}|}+\varepsilon\right)\,|C_{2}|.$
Equivalently, at most $\mu-1$ columns of $C_{2}$ are discarded.
(Completeness statement). Then for every block $i\in[q_{2}]$
[…]
In particular, the set of genuine rows of every embedded $\mathcal{M}_{1}^{(i)}$ is fully contained in $\mathcal{M}_{2}[R_{i},C^{\prime}]$; no authentic row is lost. We therefore say that each $\mathcal{M}_{1}$ inside $\mathcal{M}_{2}$ is complete.
Proof of Lemma 22.
Fix an outer index $i\in[q_{2}]$ and abbreviate
$C_{2}^{(i,c)}\;:=\;\{\mathbf{d}\in C_{2}\mid d_{i}=c\},\qquad c\in C_{1}.$
1. A multiplicity lower bound.
Because $C_{2}=S_{q_{2},t_{2}}(C_{1})$ is built with accuracy parameter $\varepsilon=\varepsilon_{q_{2},t_{2}}$, every pattern in any coordinate is $\varepsilon$–close to uniform. Hence for every $c\in C_{1}$ (and for every $i\in[q_{2}]$)
$|C_{2}^{(i,c)}|\;\;\geq\;\;\left(\tfrac{1}{|C_{1}|}-\varepsilon\right)\,|C_{2}|\;=:\;\mu.$
2. Choosing the fullness threshold.
Let $C^{\prime}\subseteq C_{2}$ be such that
$|C^{\prime}|\;\;\geq\;\;|C_{2}|-\mu\;=\;\left(1-\tfrac{1}{|C_{1}|}+\varepsilon\right)\,|C_{2}|.$
Equivalently, fewer than $\mu$ columns of $C_{2}$ are deleted.
3. No column of $\boldsymbol{C_{1}}$ disappears.
Take any $c\in C_{1}$. From ($\star$) there are at least $\mu$ columns in $C_{2}$ whose $i$-th coordinate equals $c$. Since at most $\mu-1$ columns are dropped when forming $C^{\prime}$, at least one column survives with $d_{i}=c$. Thus $C^{\prime}\cap C_{2}^{(i,c)}\neq\varnothing$ for every $c\in C_{1}$.
[…]
Row $(i,c)$ of $\mathcal{M}_{2}$ depends only on the $i$-th coordinate of the column. Because for every $c\in C_{1}$ some column with $d_{i}=c$ remains in $C^{\prime}$, the entire set of distinct rows $\{(i,c)\mid c\in C_{1}\}$ is contained in $\mathcal{M}_{2}[R_{i},C^{\prime}]$. What may be missing are duplicates: if several columns share the same pair $(i,c)$, discarding a subset of them deletes only redundant copies of that row. Duplicates are irrelevant for deterministic communication complexity, so we may identify $\mathcal{M}_{2}[R_{i},C^{\prime}]$ with $\mathcal{M}_{1}^{(i)}$ “up to the deletion of duplicated rows”.
5. Completeness.
Since the argument is uniform in $i$, every embedded $\mathcal{M}_{1}^{(i)}$ is preserved inside $\mathcal{M}_{2}[R_{i},C^{\prime}]$ in the above sense. Hence each copy of $\mathcal{M}_{1}$ inside $\mathcal{M}_{2}$ is complete, finishing the proof. ∎
Lemma 23 (Shape of a vector-augmented $\mathcal{M}_{1}$ block inside $\mathcal{M}_{4}$).
Fix an outer index $i\in[q_{2}]$ and a dimension label $c^{\star}\in C_{1}$. Consider the sub-matrix of $\mathcal{M}_{4}$ obtained by
  1. a)
    restricting the rows to
[…]
    and
  2. b)
    restricting the columns to
    $C^{\star}\;:=\;\left\{\,\left(k,(c^{\star},c^{\star},c^{\star},c^{\star})\right)\in C_{4}\right\}\;\;\subseteq\;C_{4}.$
Let $i:=|\,R^{\star}\setminus(\{i\}\times C_{1})|$ be the number of non-zero vector rows present for coordinate $c^{\star}$; thus $0\leq i\leq 4$.
After deleting the redundant duplicate rows that the relaxed interlace creates, the restricted matrix $\mathcal{M}_{4}[R^{\star},C^{\star}]$ is exactly
$\phi\left\langle\,2^{a}-1+i\,\right\rangle.$
In particular
  1. (i)
    when no vector row is present $(i=0)$ we recover the baseline copy $\phi\left\langle 2^{a}-1\right\rangle$;
  2. (ii)
    whenever at least one but not more than $2$ distinct vector rows are present, the matrix is a proper super-game of $\phi\left\langle 2^{a}\right\rangle$ and a sub-game of $\phi\left\langle 2^{a}+1\right\rangle$;
Informal proof.
Fix an outer index $i\in[q_{2}]$ and a dimension label $c^{\star}\in C_{1}$. Inside $\mathcal{M}_{4}$ the rows with index set $R_{0}:=\{i\}\times C_{1}$ already form a copy of $\mathcal{M}_{1}=\phi\left\langle 2^{a}-1\right\rangle$. Hence, before we look at any vector–rows, the restricted matrix is $\phi\left\langle 2^{a}-1\right\rangle$.
Every vector whose unique $1$ lies in coordinate $c^{\star}$ is encoded in exactly those columns whose four selector coordinates all equal $c^{\star}$. In these columns the row corresponding to that vector is one non-default row of the 5-fold interlace $\phi\left\langle 5\right\rangle$; the four remaining rows of the same interlace are all default and can be discarded without changing the set of distinct rows. Thus each such vector contributes a single extra copy of $\phi$.
Because interlacing is additive in the number of layers, appending $i$ single-row copies of $\phi$ to the original $\phi\left\langle 2^{a}-1\right\rangle$ yields
$\phi\left\langle(2^{a}-1)+i\right\rangle\;=\;\phi\left\langle 2^{a}-1+i\right\rangle,$
exactly as claimed. (Possible duplicate rows introduced by the relaxed interlace can always be removed; they play no rôle in deterministic communication complexity.)
Finally:
  1. (i)
    If $i=0,1$ we recover $\phi\left\langle 2^{a}-1\right\rangle$ or $\phi\left\langle 2^{a}\right\rangle$.
  2. (ii)
    If $1<i$ we obtain a subgame of $\phi\left\langle 2^{a}+1\right\rangle$ giving the two cases stated in the lemma.
Lemma 24.
The row set $R_{4}$ and the column set $C_{4}$ of $\mathcal{M}_{4}$ satisfy
$|R_{4}|,\;|C_{4}|\;\in\;\operatorname{poly}(d),$
i.e. their cardinalities are bounded above by some fixed polynomial in the dimension $d$ of the original $\{0,1\}$–vector bin-packing instance. Therefore, the final matrix $\mathcal{M}_{4}$ has polynomial size:
$|R_{4}\times\;C_{4}|\in\operatorname{poly}(d).$
Proof (Sketch).
Rows. $|R_{4}|=|R_{3}|+n$ where $R_{3}=[4]\times C_{2}$ and $n\leq 4d$ is the number of (padded) input vectors. It therefore suffices to upper-bound $|C_{2}|$.
Columns of $\boldsymbol{C_{2}}$. $C_{2}=S_{q_{2},t_{2}}(C_{1})$ is the AGHP reservoir with $q_{2}=\left\lceil d\right\rceil_{\!2}=d$, $|C_{1}|=\operatorname{poly}(\log d)$, and constant independence parameter $t_{2}$. The AGHP size bound $|S_{q,t}(C_{1})|\leq(q|C_{1}|)^{O(t)}\operatorname{polylog}q$ gives $|C_{2}|=\operatorname{poly}(d)$. Hence $|R_{3}|=4|C_{2}|=\operatorname{poly}(d)$ and $|R_{4}|=\operatorname{poly}(d)+O(d)=\operatorname{poly}(d)$.
Columns. $C_{4}=[32]\times R_{2}^{\,4}$ with $R_{2}=[q_{2}]\times C_{1}$. Because both $q_{2}$ and $|C_{1}|$ are polynomial in $d$, $|R_{2}|=\operatorname{poly}(d)$ and therefore $|C_{4}|=32\,|R_{2}|^{4}=\operatorname{poly}(d)$.
Thus both row and column sets of $\mathcal{M}_{4}$ are bounded by a fixed polynomial in $d$. ∎
Lemma 25.
$M_{4}$ can be constructed in polynomial time.
Proof (Sketch).
Lemma 24 already gives $|R_{4}|,|C_{4}|=\operatorname{poly}(d)$, so writing the $|R_{4}|\!\times\!|C_{4}|$ truth table of $\mathcal{M}_{4}$ takes $\operatorname{poly}(d)$ time. Each earlier stage ($\phi\!\to\mathcal{M}_{1}\!\to\mathcal{M}_{2}\!\to\mathcal{M}_{3}$) is formed by explicit formulas on matrices whose size is preserved within the same polynomial bound, so every intermediate matrix is generated in $\operatorname{poly}(d)$ time as well. Hence the whole pipeline runs in polynomial time. ∎
Theorem 3 (NP–hardness of deterministic communication complexity).
Given the truth table of a Boolean function $f\colon\{0,1\}^{n}\times\{0,1\}^{n}\to\{0,1\}$ and an integer $k$ (encoded in binary), deciding whether its deterministic communication complexity satisfies $D(f)\geq k$ is NP-hard.
Proof.
Using Lemma 23, if vectors can be partitioned so that none add up to 22 in a dimension then we end up in case (i), whereas if they cannot, we end up in case (ii). Lemma 25 ensures that the construction can be computed in polynomial time. ∎
5 Sketch for ETH-based Inapproximability
Road-map and disclaimer.
This section provides a sketch-level outline of the ETH-based additive inapproximability result announced in the abstract. We supply all definitions and state every key lemma, but we give only proof sketches whenever (i) a complete argument would duplicate material already proved in Section 3, or (ii) the underlying ideas are still being finalised and have not yet been worked out in full detail. Exact parameter settings, the explicit asymptotic form of the sub-logarithmic gap $\alpha(n)$, and a complete completeness/soundness calculation will appear in future versions (arxiv updates and conference/journal submission) once these details are fully fleshed out.
Definition 13 (Near $(\varepsilon,\delta,b)$-robust matrix).
For an integer $b\geq 0$, $\delta\in(0,\frac{1}{\sqrt{2}}-\frac{1}{2}]$ and a Boolean matrix $M$, assume that $M$ satisfies the following robustness conditions
$D\left(M\right)$   $\;\;\geq\;\;1$ (Condition 1)
$D\left(M\left[1,2^{-b},\tfrac{1}{2}+\delta\right]\right)$ $\;\;\geq\;\;D\left(M\right)-1$ (Condition 2)
$D\left(M\left[1,1-\varepsilon,\tfrac{1}{2}+\delta\right]\right)$ $\;\;\geq\;\;D\left(M\right)$ (Condition 3)
Then we call $M$ a near $(\varepsilon,\delta,b)$-robust matrix.
Lemma 26.
Given a near $(\varepsilon,\delta,b)$-robust matrix $M$, there exists $T\in\mathbb{N}$ such that for all $T\leq t\leq 2^{b}$ where $t$ is a power of $2$, and $1\leq\Delta\leq\log{t}$, we have
$D\left(M\left[\dfrac{(2^{\Delta-1}+2)t}{2^{\Delta}},2^{-b}t,2^{-\,2^{0.49\sqrt{\log t}}}\right]\right)\;\;\geq\;\;D\left(M\right)\;+\;\log t-1.$
Proof (Sketch).
The proof follows the exact same approach as Lemma 13. The only difference is that in the last step one cannot use $(\delta,b)$-robustness to substitute $D\left(M\left[1,2^{-b},\frac{1}{2}+\delta\right]\right)$ by $D\left(M\right)$ and must instead settle for $D\left(M\right)-1$ using near $(\varepsilon,\delta,b)$-robustness. ∎
Lemma 27.
Given a near $(\varepsilon,\delta,b)$-robust matrix $M$, there exists $T\in\mathbb{N}$ such that for all $T\leq t\leq 2^{b}$ where $t$ is a power of $2$, and $1\leq\Delta\leq\log{t}$, we have
$D\left(M\left[t,2^{-b}t,2^{-\,2^{0.49\sqrt{\log t}}}\right]\right)\;\;\geq\;\;D\left(M\right)\;+\;\log t.$
Proof (Sketch)..
We can inductively show that we may regain the lost bit in Lemma 26 since $t$ being a power of two means that all partitions must separate the outer indices nearly exactly or contain more than $\dfrac{(2^{\Delta-1}+2)t}{2^{\Delta}}$ interlaced functions which will hit the lower bound of Lemma 27. ∎
Theorem 4 (Additive inapproximability under ETH).
Assume the Exponential Time Hypothesis (ETH). There exists a non-decreasing function $\alpha(n)\in\omega(1)$ and, for every fixed $k\geq 2$, infinitely many input lengths $n$ such that one can construct a Boolean function
$f\;:\;\{0,1\}^{n}\to\{0,1\},$
computable by a family of $k$-alternating polynomial-time circuits, with the property that
[…]
Equivalently, under ETH no polynomial-time algorithm, given the truth table of a sub-logarithmic-alternation function, can approximate its deterministic communication complexity to within a sub-logarithmic additive gap.
Proof (Sketch) .
  1. 1.
    Base reduction. Starting from the {0,1}-Vector Bin Packing problem, we apply the structural lemmas of Section 3 to replicate the steps of Section 4 but where $\mathcal{M}_{3}$ interlaces $\Theta(\log{n})$ copies of $\mathcal{M}_{2}$. Starting from $\{0,1\}$-Vector Bin packing helps in providing robustness guarantees for the matrix obtained, which will be needed to nest the construction. This ensures that $\mathcal{M}_{4}$ is a near-robust matrix, which is defined similarly to robust matrices but where the complexity may drop by $1$ instead of remain invariant under the robustness conditions. Interlacing $\mathcal{M}_{2}$ a logarithmic number of times gives a quasi-polynomial blow-up.
  2. 2.
    Restoring robustness. Applying the interlacing operation, applied with logarithmic depth, converts $\mathcal{M}_{4}$ into a fully robust matrix, which will serve as the seed matrix $\phi$ for the next step. Again this causes quasi-polynomial blow-up. The lower-bound proof on the interlace operation has to be slightly adapted to apply to near-robust matrices, sketched in Lemma 27.
  3. 3.
    Iterated blow-up. Re-apply the base construction to $\phi$, treating it as the new seed. Each iteration increases instance size only by a quasi-polynomial factor. Because a protocol for our function needs to separate the seed matrix at each nested level of the construction, and that separation will have a difference of cost of $1$ bit depending on the satisfiability of the vector bin packing instance, each nested level amplifies the communication complexity gap by $1$. The construction can be nested $\alpha(n)\in\omega(1)$ times before becoming too large.
Precise parameter choices, the exact form of $\alpha(n)$, and a full completeness/soundness calculation will appear in the forthcoming full version. ∎
6 Discussion
Our results settle a 45-year-old question posed by Yao by proving that computing deterministic communication complexity $D\left(f\right)$ is NP-hard even for instances in which the optimal protocol uses only a constant number of alternations. The proof introduces a new relaxed interlacing gadget that keeps the instance size polynomial while preserving the projection properties of the full interlace. Because the construction is purely combinatorial, it can be transplanted into other reductions with little effort, positioning relaxed interlacing as a general-purpose tool for future lower-bound work.
6.1 Relationship to Concurrent Work
While this manuscript was being written, Hirahara et al. [3] independently showed NP-hardness of deciding $D(f)\geq k$ using an unbounded number of alternations, which they highlight as an inherent limitation of their approach. Our constant-alternation hardness resolves this issue: it rules out the possibility (assuming $P\neq NP$) that any function $f$ with bounded-round optimal protocol admits an efficient exact algorithms to compute $CC(f)$ even if more general functions do not. Note that this is distinct from the result by Hirahara et al. [2] which shows that if the protocol is only allowed to alternate a constant number of times, deciding communication complexity is NP-hard. Our reduction shows NP-hardness for functions that can be solved optimally in constant rounds.
Moreover, our ETH-based additive inapproximability work sketches out how to show that approximating $D(f)$ within $\alpha(n)\in\omega(1)$ is hard under standard fine-grained assumptions, a direction also mentioned as wide-open in the concurrent work.
6.2 Technical Novelties
  1. a)
    Relaxed interlacing. By sampling column tuples from an almost $t$-wise-independent reservoir, we circumvent the exponential blow-up of classical interlacing while having it behave similarly to the full interlace operation from the viewpoint of communication complexity. Furthermore, the self-similarity of the construction (intakes robust matrices, outputs robust matrices) enables recursive embedding, which powers the ETH gap-amplification.
  2. b)
    Block-separation lemma. The near-exact separation theorem (Theorem 2) forces every optimal protocol to spend essentially its entire budget isolating individual interlaced games. This structural control lets us encode NP-complete structure using only two extra bits of complexity and, more importantly, to iterate the construction without losing hardness.
6.3 Further Considerations
The current work builds on a framework developed in [4] and extends the tools to obtain better control the size of the constructed functions. However, the current framework still has limitations with regards to this objective that the authors strongly believe are not fundamental. Specifically, the nested construction currently suffers from the inability of the relaxed interlace operation to deal with polynomial size basis without incurring a quasi-polynomial blow-up. This is linked to it relying on a ”hard-seed” function constructed from fully interlaced functions. The present paper relaxes classical interlacing only up to $t$-wise almost-independence, which already suffices for the sub-logarithmic inapproximability gap proved above. We believe, however, that an even looser, and ultimately more powerful, variant is possible.
Towards an expander–style reservoir. Fix a constant $c>1$ and let $S\subseteq[p]^{q}$ be the multiset of column tuples used by the relaxed interlace. Instead of containing every pair $(i,j)$ whenever two coordinates are inspected (our current requirement when the independence threshold is $2$), it should suffice that $S$ satisfy the following coordinate-expander property:
For every subset $P^{\prime}\subseteq[p]$ with $|P^{\prime}|\geq|p|/c$ and for every $j\in[p]$ there exists an $i\in P^{\prime}$ with $(i,j)\in S$.
More generally, the natural $k$-ary analogue ensures that any “large” set on one coordinate expands to all values on each of the remaining k1k-1 coordinates.
Such an expander reservoir would allow for similar style structural lemmas to the ones we rely on, yet also allow us to interlace a matrix of polynomial width a polynomial number of times while keeping the overall instance size polynomial [^3] The present construction allows us to interlace a polylogarithmic width matrix a polynomial number of times whilst keeping the construction polynomial in size. (rather than quasi-polynomial). Iterating the hardness boost then yields NP-hardness of approximating $D(f)$ within a constant additive error. Constructing an explicit family of such reservoirs (e.g. via lossless condensers or high-degree bipartite expanders) and extending the block-separation analysis to this setting remains an enticing direction for future work.
Exploring the space of possible reductions. The structural theorems in Section 3 open up a whole space of possible reductions. We believe that a more thorough exploration of the possible constructions and starting problems might directly yield inapproximability results, independently of, and complementary to, nesting the construction using self-similarity.

### v2

@@
 These can be re‑used as black boxes in future lower‑bound constructions.
-Only constant alternations required.
-The optimal protocol on the instances our reduction maps onto require only constant round alternation.
-This proves NP-hardness for unrestricted rounds of alternation on a more restrictive class of functions than those with optimal protocols requiring a growing number of rounds of communication.
-This also separates our result from recent concurrent and independent work by Hirahara et al.
-[3], whose hardness proof relies on unbounded alternations.
 Recursive embedding and gap amplification.
@@
 Beyond settling the complexity of communication complexity, we expect it to be of use in the broader understanding of communication complexity.
-footnotetext: Concurrent and independent work by Hirahara, Ilango, and Loff [3] also establishes NP-hardness of deciding $D(f)\geq k$, but their reduction requires an unbounded number of alternations and does not extend to ETH-based additive inapproximability.
+footnotetext: Concurrent and independent work by Hirahara, Ilango, and Loff [2] also establishes NP-hardness of deciding $D(f)\geq k$, but their reduction does not address additive inapproximability.
 1 Preliminaries
@@
 $\delta$ baseline column margin $0<\delta\leq\frac{1}{\sqrt{2}}-\tfrac{1}{2}$ Def. 11 Fixed for whole paper
-$\varepsilon_{q,t}$ AGHP error of (q,t)(q,t)‑balanced set $\varepsilon_{q,t}:=(qt|C|)^{-c}$ §2 Drives $y^{\prime}$ in Lem. 10
-$q$ total outer blocks $q\geq t$; $q=rt$ in Lem. 14 Def. 10
-$t$ # of fully independent blocks $1\leq t\leq 2^{b}$; power of two; $\Delta\leq\log t$ Lem. 10 Lem. 13 Inner interlace factor
-$r$ replication with $q=rt$ $r\geq 1$ (int.) Lem. 14 Adds $\lceil\log r\rceil$ to budget $B$
+$\varepsilon_{q,t}$ AGHP error of (q,t)(q,t)‑balanced set $\varepsilon_{q,t}:=(qt)^{-c}$ §2 Drives $y^{\prime}$ in Lem. 10
+$q$ total outer blocks $q\geq t$; $q=rt$ in Thm. 2 Def. 10
+$t$ # of fully independent blocks $1\leq t\leq 2^{b}$; power of two; $\Delta\leq\log t$ Lem. 10,13 Inner interlace factor
+$r$ replication with $q=rt$ $r\geq 1$ (int.) Thm. 2 Adds $\lceil\log r\rceil$ to budget $B$
 $x$ row density for each $p$ interlaced functions in $\cdot\left[p,x,y\right]$ $0<x\leq 1$; often $x=2^{-b}$ or $2^{-B}$ Def. 8 Must satisfy $x\leq 2^{-k}$ in Lem. 4
@@
 $h,h^{\prime}$ column‑density maps $(1+\varepsilon)h^{\prime}(t)2^{-B}r+\varepsilon\geq h(t)$ (Extension Theorem). Otherwise monotone
-$B^{\prime}$ One round budget in Lem. 14 $B=\log t+\lceil\log r\rceil$ Lem. 14 Any optimal protocol must use all $B$ bits
+$B^{\prime}$ One round budget in Thm. 2 $B^{\prime}=\log t+\lceil\log r\rceil$ Thm. 2 Any optimal protocol must use all $B$ bits
 1.2 Definitions
@@
   1. i)
-    an integer $q\geq 1$ (number of interlacing layers),
+    an integer $q\geq 1$, (number of interlacing layers)
   2. ii)
-    a balance threshold $t\geq 1$ (number of columns we control),
+    a balance threshold $t\geq 1$, (number of columns we control)
   3. iii)
@@
 We then define a deterministic accuracy level
-$\varepsilon_{q,t}:=(qt|C|)^{-c}$
+$\varepsilon_{q,t}:=(qt)^{-c}$
 Bit-level representation.
@@
 With $k:=t\ell$, the construction of Alon et al. [1] yields a multiset [^1] A set would suffice, but the original analysis is stated for multisets. $\mathcal{B}\subseteq\{0,1\}^{q\ell}$ of size
-||= $\mathcal{B}|\;=\;2^{\,O\!\left(\log\log(q\ell)+k+\log(1/\varepsilon_{q,t})\right)}\;=\;(qp)^{\,O(t)}\cdot\operatorname{polylog}(q),$ (3)
+||= $\mathcal{B}|\;=\;2^{\,O\!\left(\log\log(q\ell)+k+\log(1/\varepsilon_{q,t}p^{t})\right)}\;=\;O(\log(q\ell))\cdot\ell^{\,O(t)}\cdot\operatorname{poly}(q\cdot p^{t}),$ (3)
 such that the uniform distribution on $\mathcal{B}$ is $\varepsilon_{q,t}$-close (in statistical distance) to $k$-wise independence.
-Importantly, this set can be constructed in $(qp)^{\,O(t)}\cdot\operatorname{polylog}(q)$ time.
+Importantly, this set can be constructed in $O(|\mathcal{B}|)$ time.
 Definition 9 (Balanced column domain).
@@
 For every index set $J\subseteq[q]$ with $|J|\leq t$ the marginal of $\mathcal{S}$ on $J$ is $\varepsilon_{q,t}$-close to the uniform distribution on $C^{J}$; equivalently, every pattern in $C^{J}$ appears with frequency
-$\frac{1}{p^{|J|}}\;\pm\;\varepsilon_{q,t}.$ (4)
+$\frac{1}{p^{|J|}}\;\pm\;\frac{\varepsilon_{q,t}}{p^{|J|}}.$ (4)
 Proof.
@@
 Remark.
-The choice $\varepsilon_{q,t}:=(qt|C|)^{-c}$ is far smaller than needed in every later bound, yet keeps $|\mathcal{S}|$ polynomial in all input parameters.
+The choice $\varepsilon_{q,t}:=(qt)^{-c}$ is far smaller than needed in every later bound, yet keeps $|\mathcal{S}|$ polynomial in all input parameters.
 Hence we dispense with a user-supplied accuracy parameter for the rest of the paper.
@@
 Only the $|\mathcal{S}|$ allowable column tuples are used, keeping the domain size
-$|[q]\times R|\cdot|\mathcal{S}|\;=\;q\,|R|\,\cdot\,2^{\,O\!\left(\log\log(q\ell)+t\ell+\log(1/\varepsilon)\right)}$
+$|[q]\times R|\cdot|\mathcal{S}|\;=\;q\,|R|\,\cdot\,2^{\,O\!\left(\log\log(q\ell)+t\ell+\log(1/\varepsilon p^{t})\right)}$
 polynomial in the natural parameters instead of exponential in $q$.
@@
 Define
-$y^{\prime}\;:=\;1-\left(1+\varepsilon\right)\,(1-y)\;=\;y-\varepsilon(1-y)\;=\;(1+\varepsilon)\,y-\varepsilon.$
+$y^{\prime}$ $=\frac{y}{1+\varepsilon}.$
 Then there exist $R^{\ast}\subseteq R^{\prime}$ and $C^{\ast}\subseteq C^{\prime}$ such that
@@
 Because $R^{\prime}$ is (Q,T)(Q,T)‑equipartitioned and $|Q|=t$, choose exactly $T$ rows from each block $\{q\}\times R$ with $q\in Q$ and discard every other row.
-Denote the surviving set by $R^{\ast}\subseteq R^{\prime}.$ Then $R^{\ast}$ is an exact (Q,T)(Q,T)‑equipartition, hence $\left|R^{\ast}\right|=tT$ and condition (i) of Definition 8 with parameter $x$ is met.
+Denote the surviving set by $R^{\ast}\subseteq R^{\prime}.$ Then $R^{\ast}$ is an exact (Q,T)(Q,T)‑equipartition, hence $\left|R^{\ast}\right|=tT$ and condition (i) of 8 with parameter $x$ is met.
 Step 2: projecting the columns.
 Let $\pi_{Q}:\mathcal{S}\to C^{t}$ be the natural projection that keeps the $t$ coordinates indexed by $Q$.
-Write $D:=\pi_{Q}(C^{\prime})\subseteq C^{t}.$ Because $C^{\prime}$ is a sub‑multiset of $\mathcal{S}$, $D$ inherits multiplicities.
-By the balance Lemma 8, the uniform distribution on $\mathcal{S}$ is $\varepsilon$‑close to tt‑wise independence; equivalently, every pattern in $C^{t}$ occurs in $\mathcal{S}$ with frequency $\frac{1}{|C|^{\,t}}\pm\varepsilon$.
-A standard averaging argument shows that at most an $\varepsilon$‑fraction of the tuples in $C^{t}$ are undersupplied compared with the ideal uniform frequency.
-Keeping only the well‑supplied patterns,
-$D^{\ast}\;:=\;\left\{\,\mathbf{c}\in C^{t}\;\big{|}\;\text{frequency of $\mathbf{c}$ in $C^{\prime}$ }\geq\left(1-\varepsilon\right)\frac{|C^{\prime}|}{|C^{t}|}\right\},$
-we lose at most an $\varepsilon$‑fraction of columns from $C^{\prime}$. More precisely,
-$|D^{\ast}|\;\geq\;|C^{t}|\;\left(y-\varepsilon(1-y)\right)\;=\;|C^{t}|\,y^{\prime}.$
-Step 3: lifting back to $\mathcal{S}$.
-For every pattern $\mathbf{c}\in D^{\ast}$ keep one representative column in $C^{\prime}$ whose projection is $\mathbf{c}$; call the resulting set $C^{\ast}\subseteq C^{\prime}$.
-By construction $|C^{\ast}|=|D^{\ast}|\geq|C^{t}|\,y^{\prime}$.
-Step 4: identifying the subgame.
+Write $D:=\pi_{Q}(C^{\prime})\subseteq C^{t}.$ Because $C^{\prime}$ is a sub‑multiset of $\mathcal{S}$, $D$ inherits multiplicities, by the balance Lemma 8, the uniform distribution on $\mathcal{S}$ is $\varepsilon$‑close to tt‑wise independence; equivalently, every pattern in $C^{t}$ occurs in $D$ with frequency $\frac{1}{|C|^{\,t}}\pm\varepsilon=\frac{1\pm\varepsilon}{|C|^{\,t}}$.
+Each pattern occurs in $\pi_{Q}(\mathcal{S})$ with frequency at most $\frac{1+\varepsilon}{|C|^{\,t}}$, $D$ contains at least contains $\frac{|C|^{t}}{1+\varepsilon}y$ unique patterns.
+More precisely, let $C^{\ast}$ be the set that contains one representative of each pattern in $D$; then
+$\frac{|C^{\ast}|}{|C|^{t}}\geq\frac{y}{1+\varepsilon}=:y^{\prime}.$
+Step 3: identifying the subgame.
 Re‑index the rows in $R^{\ast}$ as $([t]\times R)\ni(i,r)$ with $i$ the coordinate in $Q$ (keeping the original order), and re‑index each column in $C^{\ast}$ by its QQ‑projection in $C^{t}$.
-Under that relabelling the submatrix $N^{\ast}:=\varepsilon\!\left(N;R^{\ast},C^{\ast}\right)$ is exactly the classical interlace $M\langle t\rangle$ restricted to $R^{\ast}_{\text{canon}}\times D^{\ast}.$ The row set is an exact TT‑equipartition, and $|D^{\ast}|=\lceil|C|^{t}\,y^{\prime}\rceil$ by the previous step, hence $N^{\ast}\in M\left[t,x,y^{\prime}\right].$
+Under that relabelling the submatrix $N^{\ast}:=\varepsilon\!\left(N;R^{\ast},C^{\ast}\right)$ is exactly the classical interlace $M\langle t\rangle$ restricted to $R^{\ast}\times C^{\ast}.$ The row set is an exact TT‑equipartition, hence $N^{\ast}\in M\left[t,x,y^{\prime}\right].$
 Therefore $N^{\ast}$ is a member of the target bracket family and, by construction, a subgame of the original restriction $\varepsilon(N;R^{\prime},C^{\prime})$, completing the proof. ∎
@@
 $\geq\;\;$ $D\left(M\left[1,2^{-b},\tfrac{1}{2}+\delta\right]\right)-2$
-$\geq\;\;$ $1$
+$\geq\;\;$ $1.$
 Condition 4: By Lemma 6, we have
@@
 $\geq\;\;$ $D\left(M\left[1,2^{-b},\tfrac{1}{2}+\delta\right]\right)-1$
-$\geq\;\;$ $D\left(M\right)-1$
+$\geq\;\;$ $D\left(M\right)-1.$
 Condition 5: Since the communication complexity of a matrix is at least one greater than that of both submatrices obtained by a single partitioning, we consider the partition of $D\left(M\left[2,2^{-b+1},\tfrac{1}{4}+\tfrac{\delta}{2}\right]\right)$ along both rows and columns.
@@
 Where $h:\mathbb{N}\rightarrow(0,1]$. Writing $B:=D\left(M\right)+\log t+\left\lceil\log r\right\rceil$; $\varepsilon=\epsilon_{q,t}$ and assuming
-$(1+\varepsilon)h^{\prime}(t)\cdot 2^{-B}r-\varepsilon$ $\geq h(t)$ (7)
-$D\left(M\left[1,2^{-B},\left((1+\varepsilon)\,h^{\prime}(t)\,2^{-B}-\varepsilon\right)^{1/t}\right]\right)\;$  $\geq\;1,$ (8)
+$h^{\prime}(t)2^{-B}r/(1+\varepsilon)$ $\geq h(t)$ (7)
+$D\left(M\left[1,2^{-B},\left(h^{\prime}(t)\,2^{-B}/(1+\varepsilon)\right)^{1/t}\right]\right)\;$  $\geq\;1,$ (8)
+as well as
+$\forall k,c,\;D\left(M\left[1,2^{-B},\left(\frac{h^{\prime}(t)2^{-c}}{1+\varepsilon}\right)^{\frac{1}{t}}\right]\right)\geq D\left(M\right)-k$ (9)
+with $k\in[B-B^{\prime}]$ and $c\leq\log{t}+k$.
 If we denote
@@
   2. b)
-    There exists $i\in[B]$ for which $P(i)$ is true and $P(i+1)$ is false.
+    There exists $i\in[B^{\prime}-1]$ for which $P(i)$ is true and $P(i+1)$ is false.
+  3. c)
+    There exists $i\in[B^{\prime},B-1]$ for which $P(i)$ is true and $P(i+1)$ is false.
 Case a. Protocol spends its entire budget and the row–block invariant is never broken.
@@
     Apply the lemma with $t=|Q|$, $x=2^{-B}$, and the above $y$. It returns subsets $R^{\ast}\subseteq R_{i}$, $C^{\ast}\subseteq C_{i}$ and a subgame
-    $N^{\ast}\;\in\;M\left[t,2^{-B},y^{\prime}\right],\qquad y^{\prime}\;=\;(1+\varepsilon)\,h^{\prime}(t)\,2^{-B}-\varepsilon.$
+    $N^{\ast}\;\in\;M\left[t,2^{-B},y^{\prime}\right],\qquad y^{\prime}\;=\;h^{\prime}(t)\,2^{-B}/(1+\varepsilon).$
   3. Step 3:
@@
     By Lemma 2,
-    $D\left(N^{\ast}\right)\;\geq\;D\left(M\left[1,2^{-B},\left((1+\varepsilon)\,h^{\prime}(t)\,2^{-B}-\varepsilon\right)^{1/t}\right]\right).$
+    $D\left(N^{\ast}\right)\;\geq\;D\left(M\left[1,2^{-B},\left(h^{\prime}(t)\,2^{-B}/(1+\varepsilon)\right)^{1/t}\right]\right).$
   4. Step 4:
@@
     According to Equation 8,
-    $D\left(M\left[1,2^{-B},\left((1+\varepsilon)\,h^{\prime}(t)\,2^{-B}-\varepsilon\right)^{1/t}\right]\right)\;\geq\;1.$
+    $D\left(M\left[1,2^{-B},\left(h^{\prime}(t)\,2^{-B}/(1+\varepsilon)\right)^{1/t}\right]\right)\;\geq\;1.$
     Hence $D\left(N^{\ast}\right)\geq 1$.
-Case b. The row–block invariant is broken at the first step i+1i{+}1 ($1\leq i<B$).
+Case b. The row–block invariant is broken at the first step i+1i{+}1 ($1\leq i<B^{\prime}$).
 Let $\widetilde{M}_{i}=\varepsilon\left(\widehat{M};R_{i},C_{i}\right)$ be the unique rectangle that satisfies the invariant immediately before the (i+1)(i{+}1)-st bit is sent.
@@
     To break the invariant at step i+1i{+}1 the protocol must have asked at least $\left\lceil\log r\right\rceil$ row questions (each halves the number of viable blocks).
-Hence no more than $B-\left\lceil\log r\right\rceil$ bits have attacked the columns, so
-    $\frac{|C_{i}|}{|\mathsf{Cols}(\widehat{M})|}\;\geq\;h^{\prime}(t)\;2^{-(B-\lceil\log r\rceil)}\;\geq\;h^{\prime}(t)\,r\,2^{-B}\;=:\;y.$
+Hence no more than $B^{\prime}-\left\lceil\log r\right\rceil$ bits have attacked the columns, so
+    $\frac{|C_{i}|}{|\mathsf{Cols}(\widehat{M})|}\;\geq\;h^{\prime}(t)\;2^{-(B^{\prime}-\lceil\log r\rceil)}\;\geq\;h^{\prime}(t)\,r\,2^{-B^{\prime}}\;=:\;y.$
   2. Step 2:
@@
     Apply the lemma to $\widetilde{M}_{i}$ with
-    $t=|Q|,\qquad x=2^{-b},\qquad y=h^{\prime}(t)\,r\,2^{-B}.$
+    $t=|Q|,\qquad x=2^{-b},\qquad y=h^{\prime}(t)\,r\,2^{-B^{\prime}}.$
     It returns subsets $R^{\ast}\subseteq R_{i}$, $C^{\ast}\subseteq C_{i}$ and a subgame
-    $N^{\ast}\;\in\;M\left[t,2^{-b},y^{\prime}\right],\qquad y^{\prime}=(1+\varepsilon)\,h^{\prime}(t)\,r\,2^{-B}-\varepsilon.$
+    $N^{\ast}\;\in\;M\left[t,2^{-b},y^{\prime}\right],\qquad y^{\prime}=h^{\prime}(t)\,r\,2^{-B^{\prime}}/(1+\varepsilon).$
     By condition (7) in the theorem we have $y^{\prime}\geq h(t)$.
@@
 Therefore Case b is impossible, completing the proof of the Extension Theorem.
+Case c.
+The row–block invariant is broken at the first step i+1i{+}1 ($B^{\prime}\leq i<B$).
+Let $\widetilde{M}_{i}=\varepsilon\left(\widehat{M};R_{i},C_{i}\right)$ be the unique rectangle that satisfies the invariant immediately before the (i+1)(i{+}1)-st bit is sent.
+  1. Step 1:
+    Row and column parameters of $\widetilde{M}_{i}$.
+    By the invariant $R_{i}$ is a $(Q,T_{0})$-equipartition with
+    $t=|Q|,\qquad T_{0}=\left\lceil|R|\,2^{-b}\right\rceil.$
+    To break the invariant at step i+1i{+}1 the protocol must have asked at least $\left\lceil\log r\right\rceil$ row questions (each halves the number of viable blocks).
+Hence no more than $c\leq B-\left\lceil\log r\right\rceil$ bits have attacked the columns, so
+    $\frac{|C_{i}|}{|\mathsf{Cols}(\widehat{M})|}\;\geq\;h^{\prime}(t)\;2^{-c}$
+  2. Step 2:
+    Relaxed $\rightarrow$ classical interlace (Lemma 10). Apply the lemma to $\widetilde{M}_{i}$ with
+    $t=|Q|,\qquad x=2^{-B},\qquad y=h^{\prime}(t)\,2^{-c}.$
+    It returns subsets $R^{\ast}\subseteq R_{i}$, $C^{\ast}\subseteq C_{i}$ and a subgame
+    $N^{\ast}\;\in\;M\left[t,2^{-B},y^{\prime}\right],\qquad y^{\prime}=h^{\prime}(t)\,2^{-c}/(1+\varepsilon).$
+  3. Step 3:
+    Apply condition 9.
+    Invoking condition 9 with $k=i-B^{\prime}$ gives
+    $D\left(N^{\ast}\right)$ $\geq D\left(M\left[t,2^{-B},h^{\prime}(t)2^{-c}\right]\right)$
+    $\geq D\left(M\left[1,2^{-B},\left(h^{\prime}(t)2^{-c}\right)^{\frac{1}{t}}\right]\right)$
+    $\geq D\left(M\right)-k.$
+Implying that case c leads to complexity greater or equalt to the budget. ∎
 3.2 Near-Exact Separation of Interlaced Functions
@@
 The requirements for this are slightly stronger than what was required for our complexity bounds, but Theorem 2 gives us the control on the protocol that we need to encode reductions.
-Lemma 14.
-Fix integers $b\geq 1$, $r\geq 1$ and let $t\leq 2^{b}$ be a power of two. Choose $0<\delta\leq\tfrac{1}{\sqrt{2}}-\tfrac{1}{2}.$
-Assume $M$ is $(\delta,b)$‑robust and that
-$D\left(M\left[2,2^{-\log{r}-\log{t}},\left((1+\varepsilon)\,2^{-3\log{r}-3\log{t}}-\varepsilon\right)^{1/t}\right]\right)\;$ $\geq\;D\left(M\right)+1,$ (9)
+Theorem 2 (Near-Exact Separation Theorem).
+Fix integers $b\geq 1$, $r\geq 1$ and let $t\leq 2^{b}$ be a power of two. Choose $0<\delta\leq\tfrac{1}{\sqrt{2}}-\tfrac{1}{2}.$ For any
+$\delta^{\ast}\geq 4\left(\tfrac{1}{2}+\delta\right)^{3}-\tfrac{1}{2},$ (10)
+assume $M$ is $(\delta,b)$‑robust and let $B^{\prime}=\log t+\left\lceil\log r\right\rceil$, and $B=B^{\prime}+D\left(M\right)$, such that
+$D\left(M\left[2,2^{-b},\left(2^{-B^{\prime}}/(1+\varepsilon)\right)^{2/t}\right]\right)\;$  $\geq\;1,$ (11)
+as well as
+$\forall k,c,\;D\left(M\left[2,2^{-B},\left(\frac{2^{-c}}{1+\varepsilon}\right)^{\frac{2}{t}}\right]\right)\geq D\left(M\right)-k$ (12)
+with $k\in[B-B^{\prime}]$ and $c\leq\log{t}+k$.
 Let $q:=rt$ and write $\widehat{M}:=M\left\langle q,S_{q,t}(\mathsf{Cols}(M))\right\rangle.$ Pick $C^{\prime}$ with
-$\dfrac{|C^{\prime}|}{|\mathsf{Cols}(\widehat{M})|}\geq\tfrac{1}{2}+\delta.$ (10)
-Put $N:=\varepsilon\left(\widehat{M};\mathsf{Rows}(M),C^{\prime}\right)$ and $B^{\prime}:=\log t+\lceil\log r\rceil$.
-For every integer $c\geq 1$, every deterministic protocol on $N$ that sends at most $B^{\prime}$ bits but spends fewer than $B^{\prime}-c$ of them on the outer block index leaves a transcript leaf rectangle $R_{c}\times C_{c}$ such that
-$D\left(\varepsilon\left(\widehat{M};R_{c},C_{c}\right)\right)\;\geq\;D\left(M\right)+1.$
-Hence any protocol attaining the tight bound $D\left(M\right)+\log t+\lceil\log r\rceil$ must devote all $B^{\prime}$ block bits to isolating the $q=rt$ interlaced games.
+$\dfrac{|C^{\prime}|}{|\mathsf{Cols}(\widehat{M})|}\geq\tfrac{1}{2}+\delta^{\ast}.$ (13)
+Let $N:=\varepsilon\left(\widehat{M};\mathsf{Rows}(M),C^{\prime}\right)$, and $c$ be an integer such that every deterministic protocol on $N$ that sends at most $B^{\prime}$ bits but spends $B^{\prime}-c$ of them on the outer block index—keep $1-2^{-b}rt$ fraction of rows of each block to the same subgame—leaves a transcript leaf rectangle $R_{c}\times C_{c}$, such that for $c\geq 1$ we have
+$D\left(\varepsilon\left(\widehat{M};R_{c},C_{c}\right)\right)\;\geq\;D\left(M\right)+1,$
+and for $c=0$, the protocol must isolate all outer indices nearly exactly, we have that $R_{0}$ is a $(Q,|\mathsf{Rows}(N)|(1-2^{-b}rt))$-equipartition for some $Q\subseteq[rt]$ and $|Q|=1$.
 Proof.
-Let $B^{\prime}=\log t+\lceil\log r\rceil$ and fix an integer $1\leq c\leq B^{\prime}$. Write $s:=B^{\prime}-c$ for the number of questions the protocol does ask about the outer block index.
-Step 1: shape of the rectangle after $s$ block bits.
+Let $B^{\prime}=\log t+\lceil\log r\rceil$ and suppose $1\leq c\leq B^{\prime}$. Write $s:=B^{\prime}-c$ for the number of questions the protocol does ask about the outer block index.
+Dynamic invariant.
+Let $\mathsf{P}(i)$ be the statement: “After $i$ bits have been sent during the execution of $\Pi$ there exists a rectangle $R_{i}\times C_{i}$ in the current protocol partition for which $R_{i}$ is a $(Q,2^{-b})$-equipartition with $|Q|\geq t$.”
+We may then distinguish two cases:
+  1. a)
+    $\Pi$ exhausts its budget $B^{\prime}$ without ever breaking the invariant.
+  2. b)
+    There exists $i\in[B^{\prime}-1]$ for which $P(i)$ is true and $P(i+1)$ is false.
+  3. c)
+    There exists $i\in[B^{\prime},B-1]$ for which $P(i)$ is true and $P(i+1)$ is false.
+Case a. Protocol spends its entire budget and the row–block invariant is never broken.
+Let $\widetilde{M}=\varepsilon\left(\widehat{M};R_{B^{\prime}},C_{B^{\prime}}\right)$ be a rectangle that survives after the last communication bit.
+  1. Step 1:
+    Parameters of the surviving rectangle.
+    $R_{B^{\prime}}\text{ is a }(Q,T_{0})\text{-equipartition},\qquad|Q|=t,\qquad T_{0}=\left\lceil|\mathsf{Rows}(M)|\,2^{-b}\right\rceil.$
+    Because at most $B^{\prime}$ column bits were revealed,
+    $\frac{|C_{B^{\prime}}|}{|\mathsf{Cols}(\widehat{M})|}\;\geq\;2^{-B^{\prime}}$
+  2. Step 2:
+[…]
+    Apply the lemma with $t=|Q|$, $x=2^{-b}$, and the above $y$. It returns subsets $R^{\ast}\subseteq R_{B^{\prime}}$, $C^{\ast}\subseteq C_{B^{\prime}}$ and a subgame
+    $N^{\ast}\;\in\;M\left[t,2^{-b},2^{-B^{\prime}}/(1+\varepsilon)\right],$
+  3. Step 3:
+    Projection to 2 blocks.
+    By Lemma 2,
+    $D\left(N^{\ast}\right)\;\geq\;D\left(M\left[2,2^{-b},\left(2^{-B^{\prime}}/(1+\varepsilon)\right)^{2/t}\right]\right).$
+    According to Equation 11, we have $D\left(N^{\ast}\right)\geq 1$.
+Case b. The row–block invariant is broken at the first step i+1i{+}1 ($1\leq i<B^{\prime}$).
+Since the invariant is broken, there are at least $\left\lceil\log r\right\rceil$ bits spent on the outer block index, $1\leq c\leq\log t$.
+  1. Step 1:
+    Shape of the rectangle after $s$ block bits.
     The analysis separates rows and columns.
@@
 Since $q=rt$ we can lower bound it by $2^{B^{\prime}-1}+1$.
-Since $B^{\prime}-c$ bits are spent separating the outer indices, at least one element of the partition must contain $\left\lceil\frac{2^{B^{\prime}-1}+1}{2^{B^{\prime}-c}}\right\rceil$ outer indices with at least $2^{-B^{\prime}}|\mathsf{Rows}(M)|$ inner indices present.
+Since $B^{\prime}-c$ bits are spent separating the outer indices, each bit failed to separate outer indices would introduce at least one more interlace for one subgame, so at least one element of the partition must contain $\left\lceil\frac{2^{B^{\prime}-1}+1}{2^{B^{\prime}}-c}\right\rceil$ outer indices with at least $2^{-B^{\prime}}|\mathsf{Rows}(M)|$ inner indices present.
 Since $c\leq B^{\prime}$ this simplifies to $2^{c-1}+1$ outer indices that must be contained.
@@
     Column side. Each of the $c$ ”wasted bits” halve the current column set, so the surviving columns satisfy
-$|C_{c}|\;\geq\;\frac{|C^{\prime}|}{2^{c}}\;\geq\;\frac{(\tfrac{1}{2}+\delta)\,|\mathsf{Cols}(\widehat{M})|}{2^{\,c}},$ (11)
+    $|C_{c}|\;\geq\;\frac{|C^{\prime}|}{2^{c}}\;\geq\;\frac{(\tfrac{1}{2}+\delta^{\ast})\,|\mathsf{Cols}(\widehat{M})|}{2^{\,c}},$ (14)
     where the last inequality uses condition (13).
-Extraction. If the protocol does not separate all outer ranks, we obtain rectangle $R_{c}\times C_{c}$. Denoting $g(c,t):=\min{\left(2^{c}-1+1,t\right)}$, we may use Lemma 10 to infer
-$N_{c}:=\varepsilon\left(\widehat{M};R_{c},C_{c}\right)\;\in\;M\left[g(c,t),\,2^{-B^{\prime}}\,,\displaystyle y_{c}:=\frac{\frac{1}{2}+\delta}{2^{\,c}}\right].$ (12)
-Step 2: project onto two blocks (Condition 5).
-Apply the (Extended Maximum Projection Lemma). with $l=2$:
-$N_{c}^{\prime}\in M\left[2,2^{-\log{r}-\log{t}},\left(y_{c}\right)^{\,\frac{2}{{g(c,t)}}}\right].$
-$\left(\dfrac{\frac{1}{2}+\delta}{2^{c}}\right)^{\frac{2}{g(c,t)}}$ $=\left(\frac{1}{2}+\delta\right)^{\frac{2}{g(c,t)}}\cdot 2^{-\frac{2c}{g(c,t)}}$
-Given that $\frac{2}{g(c,t)}$ tends to zero the first term in the product goes to $1$ whilst
-$2^{\frac{-2c}{g(c,t)}}\geq\left(2^{-2\log{r}-2\log{t}}\right)^{\frac{1}{t}}\geq\left((1+\varepsilon)\,2^{-3\log{r}-3\log{t}}-\varepsilon\right)^{\frac{1}{t}}.$
-Therefore, condition 9 implies
-$D\left(N_{c}^{\prime}\right)\;\geq\;D\left(M\right)+1.$
-Step 3: monotonicity back to the protocol leaf. Because $N_{c}^{\prime}$ is a submatrix of the leaf rectangle $\widetilde{R}\times\widetilde{C}$, the Monotonicity Lemma 1 gives
-$D\left(\varepsilon\left(\widehat{M};R_{c},C_{c}\right)\right)\;\geq\;D\left(N_{c}^{\prime}\right)\;\geq\;D\left(M\right)+1.$
-This gives our gap. ∎
-Theorem 2 (Near-Exact Separation Theorem).
-Given a $(\delta,b)$-robust matrix $M$ such that
-$D\left(M\left[2,2^{-\log{r}-\log{t}},\left((1+\varepsilon)\,2^{-\log{r}-\log{t}}-\varepsilon\right)^{1/t}\right]\right)\;$ $\geq\;D\left(M\right)+1,$ (13)
-[…]
-$N\in\varepsilon\left(\widehat{M};R,C\right).$
-Then after $rt$ bits the elements of the partition must all be of the form $R^{\prime}\times\mathsf{Cols}(N)$ where $R^{\prime}\subseteq\mathsf{Rows}(N)$ is a $(Q,|\mathsf{Rows}(M)|(1-2^{-b}rt))$-equipartition for some $Q\subseteq[rt]$ and $|Q|=1$.
-Proof.
-Using Lemma 14, we may assume that the number of ”wasted bits” $c$ is 0, and hence no bit is spent partitioning the columns.
-From Condition 5, it we cannot have an element with $2^{-b}|\mathsf{Rows}(N)|$ rows from $2$ different outer indices.
-Since the number of outer indices equals the number of protocol leaves after rtr\,t bits, for any given outer index $i$ there is exactly one leaf $L_{i}=R_{i}\times C$ containing more than $|\mathsf{Rows}(N)|2^{-b}$ of the form $(i,e)$ with $e\in Rows(M)$.
-Since all rows must be mapped to a leaf, and there are rtr\,t leaves, this implies that for all $i$, $L_{i}=R_{i}\times C$ where $R_{i}$ contains at least $|\mathsf{Rows}(M)|\left(1-2^{-b}rt\right)$ rows of the form $(i,e)$ with $e\in\mathsf{Rows}(M)$.
-∎
+    Extraction. If the protocol does not separate all outer ranks, we obtain rectangle $R_{c}\times C_{c}$. We may use Lemma 10 to infer
+    $N_{c}:=\varepsilon\left(\widehat{M};R_{c},C_{c}\right)\;\in\;M\left[2^{c-1}+1,\,2^{-b-2B^{\prime}}\,,\displaystyle y_{c}:=\frac{\frac{1}{2}+\delta^{\ast}}{2^{\,c}(1+\varepsilon)}\right].$ (15)
+  2. Step 2:
+    Project onto two blocks (Condition 5). Apply the (Extended Maximum Projection Lemma). with $l=2$:
+    $D\left(N_{c}\right)\geq D\left(M\left[2,2^{-B^{\prime}},\left(y_{c}\right)^{\,\frac{2}{2^{c-1}+1}}\right]\right).$
+    Let $x=\tfrac{1}{2}+\delta\in\big{[}\tfrac{1}{2},\tfrac{1}{\sqrt{2}}\big{]}$. For any integer $c\geq 1$, since $t\mapsto t^{\frac{2}{2^{c-1}+1}}$ is increasing on $(0,\infty)$,
+    $\left(\frac{\tfrac{1}{2}+\delta^{\ast}}{2^{c}}\right)^{\frac{2}{2^{c-1}+1}}\geq\left(\tfrac{1}{2}+\delta\right)^{2}\iff\frac{4x^{3}}{2^{c}}\geq x^{2^{c-1}+1}\iff 4x^{3}\geq 2^{c}x^{2^{c-1}+1}.$
+    It remains to verify the last inequality. Set $g_{c}(x):=2^{c}x^{2^{c-1}+1}$. Then for $c\geq 2$,
+    $\frac{g_{c+1}(x)}{g_{c}(x)}=2x^{2^{c-1}}\leq 2x^{2}\leq 1,$
+    so $(g_{c}(x))_{c\geq 2}$ is nonincreasing. Moreover, $\frac{g_{2}(x)}{g_{1}(x)}=2x\geq 1$. Hence $g_{c}(x)\leq g_{2}(x)=4x^{3}$ for all $c\geq 1$.
+    Therefore, Condition 5 implies
+    $D\left(N_{c}\right)\;\geq\;D\left(M\right)+1.$
+  3. Step 3:
+    Monotonicity back to the protocol leaf. Because $N_{c}^{\prime}$ is a submatrix of the leaf rectangle $\widetilde{R}\times\widetilde{C}$, the Monotonicity Lemma 1 gives
+    $D\left(\varepsilon\left(\widehat{M};R_{c},C_{c}\right)\right)\;\geq\;D\left(N_{c}\right)\;\geq\;D\left(M\right)+1.$
+    This gives our gap.
+For $c=0$, the property holds trivially.
+Case c.
+The row–block invariant is broken at the first step i+1i{+}1 ($B^{\prime}\leq i<B$).
+Let $\widetilde{M}_{i}=\varepsilon\left(\widehat{M};R_{i},C_{i}\right)$ be the unique rectangle that satisfies the invariant immediately before the (i+1)(i{+}1)-st bit is sent.
+  1. Step 1:
+    Row and column parameters of $\widetilde{M}_{i}$.
+    By the invariant $R_{i}$ is a $(Q,T_{0})$-equipartition with
+    $t=|Q|,\qquad T_{0}=\left\lceil|R|\,2^{-b}\right\rceil.$
+    To break the invariant at step i+1i{+}1 the protocol must have asked at least $\left\lceil\log r\right\rceil$ row questions (each halves the number of viable blocks).
+Hence no more than $c\leq B-\left\lceil\log r\right\rceil$ bits have attacked the columns, so
+    $\frac{|C_{i}|}{|\mathsf{Cols}(\widehat{M})|}\;\geq\;2^{-c}$
+  2. Step 2:
+    Relaxed $\rightarrow$ classical interlace (Lemma 10). Apply the lemma to $\widetilde{M}_{i}$ with
+    $t=|Q|,\qquad x=2^{-B},\qquad y=2^{-c}.$
+    It returns subsets $R^{\ast}\subseteq R_{i}$, $C^{\ast}\subseteq C_{i}$ and a subgame
+    $N^{\ast}\;\in\;M\left[t,2^{-B},y^{\prime}\right],\qquad y^{\prime}=2^{-c}/(1+\varepsilon).$
+  3. Step 3:
+    Apply condition 12.
+    Invoking condition 12 with $k=i-B^{\prime}$ gives
+    $D\left(N^{\ast}\right)$ $\geq D\left(M\left[t,2^{-B},2^{-c}\right]\right)$
+    $\geq D\left(M\left[2,2^{-B},\left(2^{-c}\right)^{\frac{2}{t}}\right]\right)$
+    $\geq D\left(M\right)-k+1.$
+Implying that case c leads to complexity greater than the budget.
 4 Reduction from $\{0,1\}$-Vector Bin Packing
@@
 4.2 Construction
-Definition 12 (Transpose).
-The transpose of an $m\times n$ Boolean matrix $M$ is the matrix $M^{\top}\in\{0,1\}^{n\times m}$ defined by $M^{\top}(i,j)=M(j,i)$ for all $i\in[n]$ and $j\in[m]$.
 High-level overview.
@@
 4.3 Correctness Proof
-Table 3 gives a table of the parameters used to instantiate the reduction.
+Table 3 gives a table of the parameters used to instantiate the reduction.11todo: 1add $d$ and comment it is power of 2, Notation for power of 2
 Parameter Value chosen in the paper Brief description / where first introduced
@@
 Lemma 17.
-Any matrix $N$ in $\mathcal{M}_{1}^{\top}\left[1,\left\lceil q_{2}\right\rceil_{\!2},\left((1+\varepsilon_{q_{2},t_{2}})\,h^{\prime}(t_{2})\,\left\lceil q_{2}\right\rceil_{\!2}-\varepsilon_{q_{2},t_{2}}\right)^{1/t_{2}}\right]$ is $(\delta,2)$-robust.
+Any matrix $N$ in $\mathcal{M}_{1}^{\top}\left[1,\left\lceil q_{2}\right\rceil_{\!2},\left(h^{\prime}(t_{2})\,\left\lceil q_{2}\right\rceil_{\!2}/(1+\varepsilon_{q_{2},t_{2}})\right)^{1/t_{2}}\right]$ is $(\delta,2)$-robust.
 Proof Sketch.
@@
 Lemma 18.
-Any matrix $N$ in $\mathcal{M}_{2}^{\top}\left[1,\left\lceil q_{3}\right\rceil_{\!2},\left((1+\varepsilon_{q_{3},t_{3}})\,h^{\prime}(t_{3})\,\left\lceil q_{3}\right\rceil_{\!2}-\varepsilon_{q_{3},t_{3}}\right)^{1/t_{3}}\right]$ is $(\delta,2)$-robust.
+Any matrix $N$ in $\mathcal{M}_{2}^{\top}\left[1,\left\lceil q_{3}\right\rceil_{\!2},\left(h^{\prime}(t_{3})\,\left\lceil q_{3}\right\rceil_{\!2}/(1+\varepsilon_{q_{3},t_{3}})\right)^{1/t_{3}}\right]$ is $(\delta,2)$-robust.
 Proof Sketch.
@@
 Proof.
-Lemma 19 gives us the main condition (9) needed to apply Theorem 2. The proof follows the same arguments as 20 with the parameters one level down. ∎
+Lemma 18 gives us the main condition (11) needed to apply Theorem 2. The proof follows the same arguments as 19 with the parameters one level down. ∎
 Lemma 21 (Completeness of the embedded $\mathcal{M}_{1}$ copies in $\mathcal{M}_{2}$).
 Let $q_{2}$ and $t_{2}$ be the parameters in the definition of
-$\mathcal{M}_{2}\;=\;\mathcal{M}_{1}^{\top}\left\langle q_{2},\,S_{q_{2},t_{2}}(C_{1})\right\rangle,\qquad R_{2}=[q_{2}]\times C_{1},\quad C_{2}:=S_{q_{2},t_{2}}(C_{1}),$
+$\mathcal{M}_{2}\;=\;\mathcal{M}_{1}^{\top}\left\langle q_{2},\,S_{q_{2},t_{2}}(R_{1})\right\rangle,\qquad R_{2}=[q_{2}]\times C_{1},\quad C_{2}:=S_{q_{2},t_{2}}(R_{1}),$
 and set $\varepsilon:=\varepsilon_{q_{2},t_{2}}$.
 For every outer index $i\in[q_{2}]$ denote the $i$-th copy of $\mathcal{M}_{1}$ inside $\mathcal{M}_{2}$ by
-$\mathcal{M}_{1}^{(i)}\;:=\;\mathcal{M}_{2}\!\left[\,\{i\}\times C_{1},\,C_{2}\right],\qquad R_{i}:=\{i\}\times C_{1}.$
-(Multiplicity lower bound). By the almost–kk–wise independence of the reservoir $S_{q_{2},t_{2}}(C_{1})$ we have, for every $c\in C_{1}$,
-$\left\{\,\mathbf{d}\in C_{2}:d_{i}=c\right\}\;\;\geq\;\;\left(\tfrac{1}{|C_{1}|}-\varepsilon\right)\,|C_{2}|\;=:\;\mu.$
+$\mathcal{M}_{1}^{(i)}\;:=\;\varepsilon\left(\mathcal{M}_{2};\{i\}\times C_{1},C_{2}\right)\qquad R_{i}:=\{i\}\times C_{1}.$
+(Multiplicity lower bound). By the almost–kk–wise independence of the reservoir $S_{q_{2},t_{2}}(C_{1})$ we have, for every $c\in R_{1}$,
+$|\left\{\mathbf{d}\in C_{2}\mid d_{i}=c\right\}|\;\;\geq\;\;\tfrac{1-\varepsilon}{|R_{1}|}\cdot|C_{2}|\;=:\;\mu.$
 (Fullness condition). Let $C^{\prime}\subseteq C_{2}$ be any column subset such that
-$|C^{\prime}|\;\;\geq\;\;|C_{2}|-\mu\;=\;\left(1-\tfrac{1}{|C_{1}|}+\varepsilon\right)\,|C_{2}|.$
+$|C^{\prime}|\;\;\geq\;\;|C_{2}|-\mu+1\;=\;\left(1-\tfrac{1-\varepsilon}{|R_{1}|}\right)\,|C_{2}|.$
 Equivalently, at most $\mu-1$ columns of $C_{2}$ are discarded.
@@
 […]
-In particular, the set of genuine rows of every embedded $\mathcal{M}_{1}^{(i)}$ is fully contained in $\mathcal{M}_{2}[R_{i},C^{\prime}]$; no authentic row is lost.
+In particular, the set of genuine rows of every embedded $\mathcal{M}_{1}^{(i)}$ is fully contained in $\varepsilon\left(\mathcal{M}_{2};R_{i},C^{\prime}\right)$; no authentic row is lost.
 We therefore say that each $\mathcal{M}_{1}$ inside $\mathcal{M}_{2}$ is complete.
@@
 Fix an outer index $i\in[q_{2}]$ and abbreviate
-$C_{2}^{(i,c)}\;:=\;\{\mathbf{d}\in C_{2}\mid d_{i}=c\},\qquad c\in C_{1}.$
-1. A multiplicity lower bound.
-Because $C_{2}=S_{q_{2},t_{2}}(C_{1})$ is built with accuracy parameter $\varepsilon=\varepsilon_{q_{2},t_{2}}$, every pattern in any coordinate is $\varepsilon$–close to uniform.
-Hence for every $c\in C_{1}$ (and for every $i\in[q_{2}]$)
-$|C_{2}^{(i,c)}|\;\;\geq\;\;\left(\tfrac{1}{|C_{1}|}-\varepsilon\right)\,|C_{2}|\;=:\;\mu.$
-2. Choosing the fullness threshold.
-Let $C^{\prime}\subseteq C_{2}$ be such that
-$|C^{\prime}|\;\;\geq\;\;|C_{2}|-\mu\;=\;\left(1-\tfrac{1}{|C_{1}|}+\varepsilon\right)\,|C_{2}|.$
-Equivalently, fewer than $\mu$ columns of $C_{2}$ are deleted.
-3. No column of $\boldsymbol{C_{1}}$ disappears.
-Take any $c\in C_{1}$.
-From ($\star$) there are at least $\mu$ columns in $C_{2}$ whose $i$-th coordinate equals $c$.
+$C_{2}^{(i,c)}\;:=\;\{\mathbf{d}\in C_{2}\mid d_{i}=c\},\qquad c\in R_{1}.$
+No column of $\boldsymbol{R_{1}}$ disappears.
+Take any $c\in R_{1}$.
+There are at least $\mu\leq C_{2}^{(i,c)}$ columns in $C_{2}$ whose $i$-th coordinate equals $c$.
 Since at most $\mu-1$ columns are dropped when forming $C^{\prime}$, at least one column survives with $d_{i}=c$.
-Thus $C^{\prime}\cap C_{2}^{(i,c)}\neq\varnothing$ for every $c\in C_{1}$.
-[…]
-Row $(i,c)$ of $\mathcal{M}_{2}$ depends only on the $i$-th coordinate of the column.
-Because for every $c\in C_{1}$ some column with $d_{i}=c$ remains in $C^{\prime}$, the entire set of distinct rows $\{(i,c)\mid c\in C_{1}\}$ is contained in $\mathcal{M}_{2}[R_{i},C^{\prime}]$.
-What may be missing are duplicates: if several columns share the same pair $(i,c)$, discarding a subset of them deletes only redundant copies of that row.
-Duplicates are irrelevant for deterministic communication complexity, so we may identify $\mathcal{M}_{2}[R_{i},C^{\prime}]$ with $\mathcal{M}_{1}^{(i)}$ “up to the deletion of duplicated rows”.
-5. Completeness.
-Since the argument is uniform in $i$, every embedded $\mathcal{M}_{1}^{(i)}$ is preserved inside $\mathcal{M}_{2}[R_{i},C^{\prime}]$ in the above sense.
+Thus $C^{\prime}\cap C_{2}^{(i,c)}\neq\varnothing$ for every $c\in R_{1}$.
+Since the argument is uniform in $i$, every embedded $\mathcal{M}_{1}^{(i)}$ is preserved inside $\varepsilon\left(\mathcal{M}_{2};R_{i},C^{\prime}\right)$ in the above sense.
 Hence each copy of $\mathcal{M}_{1}$ inside $\mathcal{M}_{2}$ is complete, finishing the proof.
@@
 While this manuscript was being written, Hirahara et al.
-[3] independently showed NP-hardness of deciding $D(f)\geq k$ using an unbounded number of alternations, which they highlight as an inherent limitation of their approach.
-Our constant-alternation hardness resolves this issue: it rules out the possibility (assuming $P\neq NP$) that any function $f$ with bounded-round optimal protocol admits an efficient exact algorithms to compute $CC(f)$ even if more general functions do not.
-Note that this is distinct from the result by Hirahara et al.
-[2] which shows that if the protocol is only allowed to alternate a constant number of times, deciding communication complexity is NP-hard.
-Our reduction shows NP-hardness for functions that can be solved optimally in constant rounds.
-Moreover, our ETH-based additive inapproximability work sketches out how to show that approximating $D(f)$ within $\alpha(n)\in\omega(1)$ is hard under standard fine-grained assumptions, a direction also mentioned as wide-open in the concurrent work.
+[2] independently showed NP-hardness of deciding $D(f)\geq k$.
+The two approaches use different techniques.
+Whilst their work gives a simpler reduction, the structural lemmas derived in this paper give us greater flexibility in the design of reductions.
+This opens up the exploration of reductions from problems where an inapproximability gap exists.
+This is highlighted in Hirahara et al.
+[2] as a limitation of their approach, and we believe our framework opens up possibilities in that direction.
+Moreover, our ETH-based additive inapproximability work sketches out how to show that approximating $D(f)$ within $\alpha(n)\in\omega(1)$ is hard under standard fine-grained assumptions.
+This is possible thanks to the self-similarity inherent in our construction, and with further refinements it should be possible to extend inapproximability to stronger conditions when restricted to constants (i.e.
+NP-hardness of constant gap approximation).
 6.2 Technical Novelties